### v16

# The Ramanujan-Petersson and Selberg conjectures for Maass forms

Abstract. We prove the Ramanujan-Petersson conjecture for Maass forms of the group $SL(2,\mathbb{Z})$, with the help of automorphic distribution theory: this is an alternative to classical automorphic function theory, in which the plane takes the place usually ascribed to the hyperbolic half-plane. The Selberg eigenvalue conjecture for Hecke’s group $\Gamma_{0}(M)$ follows as well.
Keywords: Hecke eigenforms, Radon transform, pseudodifferential analysis. AMS classification codes: 11F37,44A12
1. Introduction
The present paper is based on the use of automorphic distributions. These are tempered distributions in the plane invariant under the action of $\Gamma=SL(2,\mathbb{Z})$ by means of linear transformations of $\mathbb{R}^{2}$. We shall reserve the phrase “automorphic functions” to functions in the hyperbolic half-plane $\mathbb{H}$ invariant under the well-known action of $\Gamma$ by fractional-linear transformations. An automorphic distribution contains as much information as a pair of automorphic functions: this will be detailed at the end of Section 3.
What we wish to stress from the beginning is that the use of automorphic distributions, as opposed to automorphic functions, is indispensable in this proof. Most important, it makes it possible, thanks to the Weyl operator calculus, to associate operators to such distributions. The Hecke operator becomes easy to deal with in terms of this realization. Another basic advantage is that, in the plane, it is the square of the Euler operator that has to be used in place of the Laplacian: this makes it considerably simpler to build general functions of this operator, as will be necessary in the spectral-theoretic last section.
Given a non-trivial character $\chi$ on $\mathbb{Q}^{\times}$, such that $|\chi(\frac{m}{n})|\leq\,|mn|^{C}$ for some $C>0$ and all nonzero integers $m,n$, and a real number $\lambda$, extract from the distribution
${\mathfrak{T}}_{\chi}(x,\xi)=\pi\sum_{m,n\neq 0}\chi\left(\frac{m}{n}\right)\,e^{2i\pi mx}\delta(\xi-n)$ (1.1)
its part homogeneous of degree $1-i\lambda$, to wit the tempered distribution
${\mathfrak{N}}_{\chi,i\lambda}(x,\xi)=\frac{1}{4}\sum_{mn\neq 0}|n|^{i\lambda}\,\chi\left(\frac{m}{n}\right)|\xi|^{-1-i\lambda}\exp\left(2i\pi\frac{mnx}{\xi}\right).$ (1.2)
While generally not invariant under the action in $\mathbb{R}^{2}$, by linear changes of coordinates, of the group $\Gamma=SL(2,\mathbb{Z})$, it is so for special pairs $\chi,\lambda$. The Ramanujan-Petersson conjecture is that, when such is the case, $\chi$ is of necessity a unitary character. This is a somewhat unusual formulation of the question, but it defines the environment in which it will be solved: the actual proof will be contained for the essential in Sections 7 to 9. We shall of course show the equivalence between this problem and its more traditional version, expressed in terms of modular forms of the non-holomorphic type.
Some approaches towards the Ramanujan-Petersson conjecture for Maass forms have been made several times: bounds of the Fourier coefficients $b_{p}$ of Hecke eigenforms by expressions $p^{\alpha}+p^{-\alpha}$, with exponents improving on the way, have been obtained [2, 10, 9, 13]. These results have for the most been obtained, so far as we understand, as corollaries of implementations of the Langlands functoriality principle. In contrast, the present work (which gives the result hoped for, to wit the bound corresponding to $\alpha=0$) is mostly an analyst’s job. Besides a short section devoted to elementary algebraic calculations on powers of the Hecke operator (which could be done indifferently in the plane or the half-plane), the rest of the paper consists of estimates and spectral theory: working in the plane here is essential.
An automorphic distribution is called modular if it is moreover homogeneous of some degree, and the special case considered above defines the notion of Hecke distribution. A large linear space of automorphic functions is algebraically isomorphic to the space of automorphic distributions invariant under the symplectic Fourier transformation ${\mathcal{F}}^{\mathrm{symp}}$. The map $\Theta$ from automorphic distributions to automorphic functions at work here, to wit the one defined by the identity
$(\Theta\,{\mathfrak{S}})(z)=\,\big\langle{\mathfrak{S}}\,,\,(x,\xi)\mapsto\exp\left(-\pi\,\frac{|x-z\,\xi|^{2}}{{\mathrm{Im\,}}z}\right)\big\rangle\,$, (1.3)
originated from pseudodifferential analysis (cf. (3.8)). It transforms modular distributions into modular forms of the non-holomorphic type, and distributions of the kind ${\mathfrak{N}}_{\chi,i\lambda}$ (Hecke distributions) into Hecke-normalized Hecke eigenforms: all Hecke eigenforms can be obtained in this way.
There is also a notion of Eisenstein distribution ${\mathfrak{E}}_{\pm\nu}$ (the two distributions so denoted are the images of each other under ${\mathcal{F}}^{\mathrm{symp}}$) corresponding to the more familiar one of Eisenstein series $E_{\frac{1-\nu}{2}}$ of the non-holomorphic type. All these matters, treated in [20], will be shortly reconsidered here, for the sake of self-containedness, and the equivalence between the two formulations of the Ramanujan-Petersson conjecture will follow. In [21], we proved that a combination of automorphic distribution theory and pseudodifferential analysis led to a criterion for the validity of the Riemann hypothesis: this project was finally put to an end in [22]. The present paper depends for the most (but not only) on Hecke distributions, and the preprint just cited on Eisenstein distributions only.
In automorphic analysis, $\mathbb{R}^{2}$ and $\mathbb{H}$ play complementary roles. The quotient of the half-plane by $SL(2,\mathbb{Z})$ and the automorphic Laplacian $\Delta$ are the right objects for Hilbert space methods. On the other hand, analysis in the plane concentrates as a start on the pair of operators
$2i\pi{\mathcal{E}}=x\frac{\partial}{\partial x}+\xi\frac{\partial}{\partial\xi}+1,\qquad 2i\pi{\mathcal{E}}^{\natural}=x\frac{\partial}{\partial x}-\xi\frac{\partial}{\partial\xi},$ (1.4)
the first of which only commutes with the action of $\Gamma$: the operator $\Delta-\frac{1}{4}$ is the transfer under $\Theta$ of $\pi^{2}{\mathcal{E}}^{2}$. We define for $h\in{\mathcal{S}}(\mathbb{R}^{2})$ and $t>0$
$\left(t^{2i\pi{\mathcal{E}}}h\right)(x,\xi)=t\,h(tx,t\xi),\qquad(t^{2i\pi{\mathcal{E}}^{\natural}}h)(x,\xi)=h(tx,t^{-1}\xi)),$ (1.5)
so that $t\frac{d}{dt}\left(t^{2i\pi{\mathcal{E}}}h\right)=(2i\pi{\mathcal{E}})\,\left(t^{2i\pi{\mathcal{E}}}h\right)$ and something similar goes with the other operator. Transposing the operator $2i\pi{\mathcal{E}}$ or $2i\pi{\mathcal{E}}^{\natural}$ is the same as changing it to its negative so that, using duality, the definition just given applies as well if one replaces $h\in{\mathcal{S}}(\mathbb{R}^{2})$ by ${\mathfrak{S}}\in{\mathcal{S}}^{\prime}(\mathbb{R}^{2})$.
On the other hand, the Hecke operator $T_{p}$ familiar to number theorists becomes in the plane the operator $T_{p}^{\mathrm{dist}}=p^{-\frac{1}{2}+i\pi{\mathcal{E}}^{\natural}}+p^{\frac{1}{2}-i\pi{\mathcal{E}}^{\natural}}\sigma_{1}$, where $\sigma_{1}$ is a simple averaging operator, a special case of the operator $\sigma_{r}$ ($r=0,1,\dots$) such that
$\left(\sigma_{r}{\mathfrak{S}}\right)(x,\xi)=\frac{1}{p^{r}}\sum_{b\,{\mathrm{mod}}\,p^{r}}{\mathfrak{S}}\left(x+\frac{b\xi}{p^{r}},\,\xi\right).$ (1.6)
Simple algebraic calculations lead to the expression of the $(2N)$th power of $T_{p}^{\mathrm{dist}}$ as a linear combination, with essentially explicit coefficients, of the operators $\left(p^{-1+2i\pi{\mathcal{E}}^{\natural}}\right)^{N-\ell}\sigma_{r}$, with $0\leq\ell\leq 2N$ and $(2\ell-2N)_{+}\leq r\leq\ell$.
Setting $\Gamma_{\infty}=\{\left(\begin{smallmatrix}1&b\\ 0&1\end{smallmatrix}\right)\colon b\in\mathbb{Z}\}$, consider the measure ${\mathfrak{s}}_{1}^{1}(x,\xi)=e^{2i\pi x}\delta(\xi-1)$. The series
${\mathfrak{B}}=\frac{1}{2}\sum_{g\in\Gamma/\Gamma_{\infty}}{\mathfrak{s}}_{1}^{1}\,\circ\,g^{-1}$ (1.7)
does not converge weakly in the space ${\mathcal{S}}^{\prime}(\mathbb{R}^{2})$ of tempered distributions, but it does so in the space of continuous linear forms on the space defined as the image of ${\mathcal{S}}_{\mathrm{even}}(\mathbb{R}^{2})$ under $\pi^{2}{\mathcal{E}}^{2}$. It is of course automorphic, and it decomposes as a superposition of modular distributions as the identity
${\mathfrak{B}}=\frac{1}{4\pi}\int_{-\infty}^{\infty}\frac{1}{\zeta(i\lambda)\,\zeta(-i\lambda)}\,{\mathfrak{E}}_{i\lambda}\,d\lambda+\frac{1}{2}\sum_{\begin{array}[]{c}r,\iota\\ r\in\mathbb{Z}^{\times}\end{array}}\frac{\Gamma(-\frac{i\lambda_{r}}{2})\,\Gamma(\frac{i\lambda_{r}}{2})}{\|\,{\mathcal{N}}^{|r|,\iota}\,\|^{2}}\,{\mathfrak{N}}^{r,\iota}\,$, (1.8)
better explained in detail in Section 6: let us just observe here that no modular distribution is missing.
The automorphic distribution $(\pi{\mathcal{E}})^{2}{\mathfrak{B}}$ is invariant under the symplectic Fourier transformation, from which it follows that it contains exactly the same information as its image under $\Theta$. Now, this image is a special case of a series introduced by Selberg [14], general versions of which are referred to as Poincaré series in the more recent literature. While the two versions, the automorphic distribution and the automorphic function which is its image under $\Theta$, are mathematically equivalent, it is only the first object that could be treated, in Sections 7 and 8, by methods of analysis. We shall devote the end of Section 8 to a justification of this point of view.
From this point on, we stay entirely within the automorphic distribution environment. We shall show that, with an integer $N$ going to infinity, and given a prime $p$, the operator $\left(p^{-\frac{1}{2}+i\pi{\mathcal{E}}^{\natural}}+p^{\frac{1}{2}-i\pi{\mathcal{E}}^{\natural}}\sigma_{1}\right)^{2N}$, which acts as a scalar on each term of (1.8), does not increase the coefficients by a factor larger than $(2\delta)^{2N}$, where $\delta$ is an arbitrary number >1>1. Note that finding almost uniform estimates for the $N$th powers of $T_{p}^{\mathrm{dist}}$ makes it possible, using $N$th roots, to deal with constants on which we have no control.
This part, based on the definition (1.7) of $\mathfrak{B}$}, constitutes the central point of our approach. Using (1.6), we shall have to find appropriate bounds, in ${\mathcal{S}}^{\prime}(\mathbb{R}^{2})$, for the individual terms $\left(p^{-1+2i\pi{\mathcal{E}}^{\natural}}\right)^{N-\ell}\sigma_{r}{\mathfrak{B}}$. Attempts at proving the desired estimate proved an extremely cumbersome task, until we found that using pseudodifferential analysis, in other words considering the operator attached to $\mathfrak{B}$} under the “quantizing map” $\Psi$ (cf. (3.5) below) saved the situation. What then remained to be done was a last spectral-theoretic development, performed by the insertion under the integral or summation sign in (1.8) of some operator $\Phi_{N}(2i\pi{\mathcal{E}})$ concentrating the resulting distribution, in some sense, near any given discrete eigenvalue $i\lambda_{r}$ of the automorphic version of $2i\pi{\mathcal{E}}$.
We wish to conclude this introduction by the observation that, considering its central role both in the present paper and in our discussion [22] of the Riemann hypothesis, pseudodifferential operator theory, an absolutely fundamental tool of PDE’s, may also find a place in number theory.
2. Homogeneous and bihomogeneous functions in the plane
We fix notation: needless to say, there is nothing original here. Given $\nu\in\mathbb{C}$ and $\delta=0$ or $1$, the function $t\mapsto|t|_{\delta}^{-\nu}=|t|^{-\nu}\,({\mathrm{sign}}\,t)^{\delta}$ on $\{$$\mathbb{R}\backslash\{0\}$ is locally summable if ${\mathrm{Re\,}}\nu<1$. The identity
$|t|_{\delta}^{-\nu}=\frac{\Gamma(1-\nu)}{\Gamma(1-\nu+k)}\,\left(\frac{d}{dt}\right)^{k}|t|^{-\nu+k}_{\delta^{\prime}},\qquad\delta^{\prime}\equiv\delta+k\,\,{\mathrm{mod}}\,2,\,\,{\mathrm{Re\,}}(k-\nu)>0,$ (2.1)
to which one may add $t^{-1}=\frac{d}{dt}\,\log\,|t|$, makes it possible to extend $|t|_{\delta}^{-\nu}$ as a tempered distribution on the line, provided that $\nu\neq\delta+1,\delta+3,\dots$. Let us introduce the shorthand
$B_{\delta}(\nu)=(-i)^{\delta}\,\pi^{\nu-\frac{1}{2}}\,\frac{\Gamma(\frac{1-\nu+\delta}{2})}{\Gamma(\frac{\nu+\delta}{2})},\qquad\nu\neq\delta+1,\delta+3,\dots,$ (2.2)
noting the functional equation $B_{\delta}(\nu)B_{\delta}(1-\nu)=(-1)^{\delta}$ and the relation $B_{0}(\nu)=\frac{\zeta(\nu)}{\zeta(1-\nu)}$, equivalent to the functional equation of the zeta function. One has
$\left({\mathcal{F}}\left(|t|_{\delta}^{-\nu}\right)\right)(\tau)\colon=\int_{-\infty}^{\infty}\,|t|_{\delta}^{-\nu}\,e^{-2i\pi t\tau}\,dt=B_{\delta}(\nu)\,|\tau|_{\delta}^{\nu-1},$ (2.3)
a semi-convergent integral if $0<{\mathrm{Re\,}}\nu<1$: if this is not the case, the Fourier transformation in the space of tempered distributions makes it possible to extend the identity, provided that $\nu\neq\delta+1,\delta+3,\dots$ and $\nu\neq-\delta,-\delta-2,\dots$. As functions of $\nu$ with values in ${\mathcal{S}}^{\prime}(\mathbb{R})$, both sides of this equation are holomorphic in the domain just indicated. Recall for reference the asymptotics of the Gamma function on vertical lines [12, p.13]
$|\Gamma(x+iy)|\sim\,(2\pi)^{\frac{1}{2}}e^{-\frac{\pi}{2}|y|}\,|y|^{x-\frac{1}{2}},\qquad|y|\to\infty.$ (2.4)
In the plane $\mathbb{R}^{2}$, we shall use the pair of commuting operators, the first of which is the Euler operator,
$2i\pi{\mathcal{E}}=x\frac{\partial}{\partial x}+\xi\frac{\partial}{\partial\xi}+1,\qquad 2i\pi{\mathcal{E}}^{\natural}=x\frac{\partial}{\partial x}-\xi\frac{\partial}{\partial\xi}.$ (2.5)
The operators $\mathcal{E}$} and $\mathcal{E}$$^{\natural}$ are formally self-adjoint in $L^{2}(\mathbb{R}^{2})$, but we shall only use them in the Schwartz space ${\mathcal{S}}(\mathbb{R}^{2})$ and its dual ${\mathcal{S}}^{\prime}(\mathbb{R}^{2})$, the space of tempered distributions. Homogeneous distributions are generalized eigenfunctions of the operator $2i\pi{\mathcal{E}}$: we need to consider also bihomogeneous functions
${\mathrm{hom}}_{\rho,\nu}^{(\delta)}(x,\xi)=|x|_{\delta}^{\frac{\rho+\nu-2}{2}}|\xi|_{\delta}^{\frac{-\rho+\nu}{2}},\qquad x\neq 0,\,\xi\neq 0.$ (2.6)
They are globally even generalized eigenfunctions of the pair $(2i\pi{\mathcal{E}},\,2i\pi{\mathcal{E}}^{\natural})$ for the pair of eigenvalues $(\nu,\rho-1)$, and they make sense as distributions in the plane provided that $\rho+\nu\neq-2\delta,-2\delta-4,\dots$ and $2-\rho+\nu\neq-2\delta,-2\delta-4,\dots$.
It is useful to keep handy the formula giving the decomposition of “general” functions $h$ in $\mathbb{R}^{2}$ into generalized eigenfunctions of the operator $2i\pi{\mathcal{E}}$. Given $h\in{\mathcal{S}}(\mathbb{R}^{2})$ and $(x,\xi)\neq(0,0)$, the function $f(t)=e^{2\pi t}h\left(e^{2\pi t}x,e^{2\pi t}\xi\right)$ is integrable and, defining
$h_{i\lambda}(x,\xi)=\widehat{f}(-\lambda)=\frac{1}{2\pi}\int_{0}^{\infty}\theta^{i\lambda}h(\theta x,\theta\xi)\,d\theta,$ (2.7)
[…]
The notation (1.5) may be regarded as justified by the equation $t\frac{d}{dt}_{\big|t=1}\left(t^{2i\pi{\mathcal{E}}}h\right)=(2i\pi{\mathcal{E}})\,h$ that follows. It is in an $L^{2}(\mathbb{R}^{2})$-setting, however, that $t^{2i\pi{\mathcal{E}}}$ takes its full sense. As defined on ${\mathcal{S}}(\mathbb{R}^{2})$, the symmetric (i.e., formally self-adjoint) operator $2\pi{\mathcal{E}}$ is essentially self-adjoint (which means that it has a unique self-adjoint extension, also denoted as $2\pi{\mathcal{E}}$): then, the map $t\mapsto t^{2i\pi{\mathcal{E}}}$ as defined in (1.5) is a one-parameter group of unitary operators, the infinitesimal generator of which, in the sense of Stone’s theorem (for instance, [23]) is the operator $2\pi{\mathcal{E}}$.
3. From the plane to the half-plane
Denote as $\mathbb{H}$ the hyperbolic half-plane $\{z=x+iy\in\mathbb{C}\colon y>0\}$, provided with its usual measure $y^{-2}dx\,dy$, invariant under the action of the group $SL(2,\mathbb{R})$ by fractional-linear transformations $\left(\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right),\,z\right)\mapsto\frac{az+b}{cz+d}$. It is a Euclidean domain, with the Laplacian $\Delta=(z-\overline{z})^{2}\frac{\partial^{2}}{\partial\,z\partial\,\overline{z}}=-y^{2}\left(\frac{\partial^{2}}{\partial x^{2}}+\frac{\partial^{2}}{\partial y^{2}}\right)$.
If ${\mathfrak{S}}={\mathfrak{S}}(x,\xi)$ (the notation $(x,\xi)$ for points of $\mathbb{R}^{2}$ is traditional in pseudodifferential analysis, from which all the new notions in this paper originated), we denote as $\Theta\,{\mathfrak{S}}$ the analytic function in $\mathbb{H}$ defined as
$\left(\Theta\,{\mathfrak{S}}\right)(z)=\,\big\langle{\mathfrak{S}},\,(x,\xi)\mapsto\exp\left(-\pi\frac{|x-z\xi|^{2}}{{\mathrm{Im\,}}z}\right)\big\rangle\,.$ (3.1)
It is immediate that the map $\Theta$ intertwines two actions of the group $SL(2,\mathbb{R})$: the one, by fractional-linear transformations in $\mathbb{H}$, and the action defined as $\left(\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right),\,\left(\begin{smallmatrix}x\\ \xi\end{smallmatrix}\right)\right)\mapsto\left(\begin{smallmatrix}ax+b\xi\\ cx+d\xi\end{smallmatrix}\right)$ in the plane. Note that in expressions $\langle{\mathfrak{S}},h\rangle$ such as the one on the right-hand side of (3.1), the use of straight brackets indicates that we are dealing with the bilinear form defined in terms of the integral. One has the transfer identity
$\Theta\left(\pi^{2}{\mathcal{E}}^{2}{\mathfrak{S}}\right)=\left(\Delta-\frac{1}{4}\right)\Theta\,{\mathfrak{S}}.$ (3.2)
Setting $\rho=\frac{|x-z\,\xi|^{2}}{{\mathrm{Im\,}}z}$, this identity can indeed be written as
$\left(\rho\,\frac{d}{d\rho}+\frac{1}{2}\right)^{2}f(\rho)=\left(\Delta-\frac{1}{4}\right)f(\rho)$ (3.3)
for every $C^{2}$ function $f$: the calculation of the right-hand side presents no complication since, taking advantage of the invariance of both operators involved under the appropriate actions of $SL(2,\mathbb{R})$, one may assume that $(x,\xi)=(1,0)$, so that $\rho=({\mathrm{Im\,}}z)^{-1}$.
The spaces $\mathbb{R}^{2}$ and $\mathbb{H}$ are two homogeneous spaces of the group $SL(2,\mathbb{R})$ and the map $\Theta$ is an associate of the dual Radon transformation, the most elementary case of Helgason’s theory [5]. More precisely [19, Prop. 2.1.1], it is obtained from the dual Radon transformation $V^{*}$ by inserting on the right-hand side the operator $\pi^{\frac{1}{2}-i\pi{\mathcal{E}}}\Gamma\left(\frac{1}{2}+i\pi{\mathcal{E}}\right)$. In view of the decreasing (2.4) of the Gamma factor at infinity on vertical lines, this is already an indication of the fact that analysis in the plane is bound to give precise results more easily than its image in $\mathbb{H}$, as obtained under $\Theta$. Actually, the operator $\Theta$ was introduced for the necessities of pseudodifferential analysis (cf. (3.8) below).
In the plane, we always use the symplectic Fourier transformation, as defined by the equation
$\left({\mathcal{F}}^{\mathrm{symp}}\,{\mathfrak{S}}\right)(x,\xi)=\int_{\mathbb{R}^{2}}{\mathfrak{S}}(y,\eta)\,e^{2i\pi(x\eta-y\xi)}dy\,d\eta,$ (3.4)
which is invariant under the action by linear changes of coordinates of $SL(2,\mathbb{R})$. One has for every tempered distribution $\mathfrak{S}$} the identity $\Theta{\mathcal{F}}^{\mathrm{symp}}\,{\mathfrak{S}}=\Theta\,{\mathfrak{S}}$, which shows that the map $\Theta$ is not quite one-to-one. But if $\mathfrak{S}$} is globally even (which will be the case considered, most of the time), it is characterized by the pair of its images under $\Theta,\,\Theta\,(2i\pi{\mathcal{E}})$: some pseudodifferential analysis is required to prove this. Introduce it as the linear map $\Psi$ (one-to-one and onto, denoted as ${\mathrm{Op}}_{2}$ in [21]) from ${\mathcal{S}}^{\prime}(\mathbb{R}^{2})$ to the space of continuous linear operators from ${\mathcal{S}}(\mathbb{R})$ to ${\mathcal{S}}^{\prime}(\mathbb{R})$ defined by the formula
$\left(v\,\big|\,\Psi({\mathfrak{S}})\,u\right)\colon=\,\big\langle\overline{v},\,\Psi({\mathfrak{S}})\,u\big\rangle\,=\,\big\langle{\mathfrak{S}},\,{\mathrm{Wig}}(v,u)\big\rangle\,,\qquad u,v\in{\mathcal{S}}(\mathbb{R}),$ (3.5)
with
${\mathrm{Wig}}(v,u)(x,\xi)=\int_{-\infty}^{\infty}\overline{v}(x+t)\,u(x-t)\,e^{2i\pi t\xi}dt.$ (3.6)
If $\mathfrak{S}$} is a globally even distribution, the operator $\Psi({\mathfrak{S}})$ preserve the parity of functions. Also, one has the general identity $\Psi\left({\mathcal{F}}^{\mathrm{symp}}\,{\mathfrak{S}}\right)u=\Psi({\mathfrak{S}})\,\overset{\vee}{u}$, with $\overset{\vee}{u}(x)=u(-x)$.
One remarks (this is no accident) that the function which enters the definition (3.1) of $\Theta$ is a composite object: indeed, taking
$u_{z}(x)=\left({\mathrm{Im\,}}(-z^{-1})\right)^{\frac{1}{4}}\,\exp\left(\frac{i\pi x^{2}}{2\overline{z}}\right),$ (3.7)
one has the identity
${\mathrm{Wig}}(u_{z},u_{z})(x,\xi)=\exp\left(-\pi\,\frac{|x-z\xi|^{2}}{{\mathrm{Im\,}}z}\right).$ (3.8)
It follows that, if $\Theta{\mathfrak{S}}=0$, one has first $\left(u_{z}\,\big|\,\Psi({\mathfrak{S}})\,u_{z}\right)=0$ for every $z\in\mathbb{H}$, then (regarding the exponent in (3.8) as the restriction to the diagonal $w=z$ of the function $\frac{2i\,(x-\overline{w}\,\xi)(x-z\,\xi)}{z-\overline{w}}$ and using a “sesquiholomorphic” argument) that $\left(u_{w}\,\big|\,\Psi({\mathfrak{S}})\,u_{z}\right)=0$ for every pair $w,z$ of points of $\mathbb{H}$. As linear combinations of functions $u_{z}$ build up to a dense subspace of ${\mathcal{S}}_{\mathrm{even}}(\mathbb{R})$, it follows that the restriction of $\Psi({\mathfrak{S}})$ to ${\mathcal{S}}_{\mathrm{even}}$ is zero. Replacing $u_{z}(x)$ by $x\,u_{z}(x)$, one shows in the same way that the restriction of $\Psi({\mathfrak{S}})$ to ${\mathcal{S}}_{\mathrm{odd}}$ is characterized by the function $\Theta\,(2i\pi{\mathcal{E}})\,{\mathfrak{S}}$.
This implies that two functions in $\mathbb{H}$, to wit $\Theta\,{\mathfrak{S}}$ and $\Theta\,(2i\pi{\mathcal{E}})\,{\mathfrak{S}}$, are needed to characterize a globally even tempered distribution $\mathfrak{S}$} in $\mathbb{R}^{2}$, and that they suffice. Pseudodifferential operator theory in an arithmetic environment, to be precise the analysis of operators $\Psi({\mathfrak{S}})$ in which $\mathfrak{S}$} is an automorphic distribution, is the central element in the analysis [22] of the Riemann hypothesis.
In the next section, we shall reconsider the Radon transform or the associated operator $\Theta$ in an arithmetic environment, more specifically in the $\Gamma$-invariant situation, with $\Gamma=SL(2,\mathbb{Z})$. We explain now some of the advantages of working in the plane. These concern analysis, not algebraic facts which, for instance the elementary calculations in Section 5, could be carried just as well in the hyperbolic half-plane. Also, there is a nice Hilbert space structure on $\)L^{2}(\Gamma\backslash\mathbb{H}$): though there exists a Hilbert space $L^{2}(\Gamma\backslash\mathbb{R}^{2})$ [19, Section 5], its definition is far from being as simple as that of $\)L^{2}(\Gamma\backslash\mathbb{H}$), because of the lack of a fundamental domain for the action of $\Gamma$ in $\mathbb{R}^{2}$. We shall go back to modular form theory when, as in Section 6, Hilbert space facts are needed.
One of the main benefits of modular distribution theory lies in the role of the Euler operator $\pi{\mathcal{E}}$: it commutes with the action of $SL(2,\mathbb{Z})$ (or $SL(2,\mathbb{R})$), an essential point in the last part (spectral localization) of this paper. Its square transfers under $\Theta$ to $\Delta-\frac{1}{4}$: but the operator $\pi{\mathcal{E}}$ does not transfer to $\mathbb{H}$. Note that the transfer under $\Theta$ of the operator $2i\pi{\mathcal{E}}^{\natural}$ is simpler, since one has the general formula $\left[\Theta\left(p^{2i\pi{\mathcal{E}}^{\natural}}{\mathfrak{S}}\right)\right](z)=[\Theta{\mathfrak{S}}](p^{2}z)$.
It is of course easier to use first-order operators, to start with when integrations by parts are needed, as will be the case in Section 7. Spectral-theoretic reasons also show the benefit of automorphic distribution theory. To find bounds for the eigenvalues of a high power of the Hecke operator $T_{p}^{\mathrm{dist}}$ (the superscript indicates that we have transferred $T_{p}$ to the plane), one lets this operator act on some distribution $\mathfrak{B}$} the decomposition of which into homogeneous components contains all Hecke distributions (cf. next section and Section 6). But, having found a satisfactory bound for the action of $\left[T_{p}^{\mathrm{dist}}\right]^{2N}$ on $\mathfrak{B}$}, one still needs to localize (spectrally) the estimate obtained to see the effect of $\left[T_{p}^{\mathrm{dist}}\right]^{2N}$ on individual Hecke distributions: this is done with the help of a well-chosen function of the operator $\mathcal{E}$}. Needless to say, functions of $\mathcal{E}$}, superpositions of the rescaling operators $t^{2i\pi{\mathcal{E}}}$, are easier to define and to use than functions of $\Delta-\frac{1}{4}$, the integral kernels of which require the use of Legendre functions in the pair invariant $\cosh d(z,w)$ on $\mathbb{H}$. The heart of the matter is that while the subgroup $SL(2,\mathbb{Z})$ of $SL(2,\mathbb{R})$ defines modular objects in either case, introducing the subgroup $SO(2)$, as needed to define the hyperbolic half-plane, can sometimes only complicate the picture.
But the use, in the present paper, of automorphic distribution theory, has a more fundamental aspect, to be made clear in Remark 8.2 at the end of Section 8. Using only automorphic function theory (in the half-plane) would not enable us to complete the proof of Proposition 8.2.
To complete the list of advantages of the automorphic distribution point of view, let us mention, though this will not enter the present paper, that the use of the Weyl calculus endows, up to a point, the space of automorphic distributions with a structure of (non-commutative) operator-algebra [20]. As mentioned above, the main advantage of piecing together automorphic distribution theory and the Weyl calculus lies in that it leads to a disproof [22] of the Riemann hypothesis.
The Radon transformation from functions on $\mathbb{H}$ to even functions on $\mathbb{R}^{2}$ is the simplest case of Helgason’s theory [5], to wit for $G=SL(2,\mathbb{R})$, in which case, with classical notation, $\mathbb{H}\sim G/K$ and the space of even functions on $\mathbb{R}^{2}$ identifies with the space of functions on $G/MN$. The link between automorphic function and automorphic distribution theory, analyzed in the next section, is the $SL(2,\mathbb{Z})$-invariant version of the same. Actually, our construction of automorphic distribution theory was made [17, section 18] under the influence of the correspondence in the Lax-Phillips scattering theory [11] between a Cauchy problem for the wave equation in $3$-dimensional spacetime with data on one sheet of a two-sheeted hyperboloid, and a Goursat (totally characteristic) problem with datum on the boundary of the light-cone. A summary of the construction can be found as Section 12 in [22]
4. Modular distributions and modular forms
This section is a summary, with minor modifications, of [20, Section 2.1] and [21, Section 6.3]. Its main point is to show that, when we have proved the Ramanujan-Petersson conjecture as formulated in the beginning of the introduction, it will entail the validity of the conjecture in its traditional formulation, involving modular forms. We first quote [20, Theor.1.2.2].
Proposition 4.1.
Given a non-trivial character $\chi$ on $\mathbb{Q}^{\times}$, such that $|\chi(\frac{m}{n})|\leq|mn|^{C}$ for some $C>0$, and $\lambda\in\mathbb{R}$, the even distribution ${\mathfrak{N}}_{\chi,i\lambda}\in{\mathcal{S}}^{\prime}(\mathbb{R}^{2})$ defined by the equation
$\big\langle{\mathfrak{N}}_{\chi,i\lambda}\,,\,h\big\rangle\,=\frac{1}{4}\,\sum_{m,\,n\neq 0}\,\chi\left(\frac{m}{n}\right)\,\int_{-\infty}^{\infty}|t|^{-1-i\lambda}\,\left({\mathcal{F}}_{1}^{-1}h\right)\left(\frac{m}{t},\,nt\right)\,dt\,,\qquad h\in{\mathcal{S}}(\mathbb{R}^{2}),$ (4.1)
satisfies the identity $\big\langle{\mathfrak{N}}_{\chi,i\lambda}\,,\,h\,\circ\,\left(\begin{smallmatrix}1&1\\ 0&1\end{smallmatrix}\right)\big\rangle\,=\,\big\langle{\mathfrak{N}}_{\chi,i\lambda}\,,\,h\big\rangle\,$ for every function $h\in{\mathcal{S}}(\mathbb{R}^{2})$. Also, it is homogeneous of degree $1-i\lambda$. Set $\chi(-1)=(-1)^{\delta}$ with $\delta=0$ or $1$, and define
$\psi_{1}(s)=\sum_{m\geq 1}\chi(m)\,m^{-s}=\prod_{p}\left(1-\chi(p)\,p^{-s}\right)^{-1},\qquad\psi_{2}(s)=\sum_{n\geq 1}(\chi(n))^{-1}n^{-s},$ (4.2)
two convergent series for ${\mathrm{Re\,}}s$ large enough. Also, set
$L(s,{\mathfrak{N}}_{\chi,i\lambda})$ $=\psi_{1}\left(s+\frac{i\lambda}{2}\right)\psi_{2}\left(s-\frac{i\lambda}{2}\right),$
$L^{\natural}(s,{\mathfrak{N}}_{\chi,i\lambda})$ $=\frac{1}{2}\,B_{\delta}\left(\frac{2-i\lambda}{2}-s\right)\,L(s,{\mathfrak{N}}_{\chi,i\lambda}),$ (4.3)
and assume that the function $s\mapsto L(s,{\mathfrak{N}}_{\chi,i\lambda})$ extends as an entire function of $s$, polynomially bounded in vertical strips. Then, the distribution ${\mathfrak{N}}_{\chi,i\lambda}$ admits a decomposition into bihomogeneous components, given as
${\mathfrak{N}}_{\chi,i\lambda}=\frac{1}{4i\pi}\,\int_{{\mathrm{Re\,}}\rho=1}L^{\natural}\left(\frac{2-\rho}{2},{\mathfrak{N}}_{\chi,i\lambda}\right)\,\,{\mathrm{hom}}^{(\delta)}_{\rho,-i\lambda}\,d\rho.$ (4.4)
It is $\Gamma$-invariant, i.e., a modular distribution, if and only if the functional equation
$L^{\natural}(s,{\mathfrak{N}}_{\chi,i\lambda})=(-1)^{\delta}\,L^{\natural}(1-s,{\mathfrak{N}}_{\chi,i\lambda})$ (4.5)
is satisfied.
Given an automorphic distribution $\mathfrak{S}$}, we set, for $k=1,2,\dots$,
$\big\langle T_{k}^{\mathrm{dist}}{\mathfrak{S}}\,,\,h\big\rangle\,=k^{-\frac{1}{2}}\sum_{\begin{array}[]{c}ad=k,\,d>0\\ b\,{\mathrm{mod}}\,d\end{array}}<{\mathfrak{S}}\,,\,(x,\,\xi)\mapsto h\left(\frac{dx-b\xi}{\sqrt{k}},\,\frac{a\xi}{\sqrt{k}}\right)>$ (4.6)
and
$\big\langle T_{-1}^{\mathrm{dist}}{\mathfrak{S}}\,,\,h\big\rangle\,=\,\big\langle{\mathfrak{S}}\,,\,(x,\xi)\mapsto h(-x,\,\xi)\big\rangle\,.$ (4.7)
As a straightforward computation shows, under $\Theta$, the operator $T_{k}^{\mathrm{dist}}$ transfers if $k\geq 1$ to the operator $T_{k}$ such that [6, p.127]
$(T_{k}f)(z)=k^{-\frac{1}{2}}\sum_{\begin{array}[]{c}ad=k,\,d>0\\ b\,\,{\mathrm{mod}}\,d\end{array}}f\left(\frac{az+b}{d}\right),$ (4.8)
to be completed by $(T_{-1}f)(z)=f(-\overline{z})$. This is the familiar collection of Hecke operators known to practitioners of (non-holomorphic) modular form theory, and it is well-known that it constitutes a commutative family of operators, generated by $T_{-1}$ and the operators $T_{p}$ with $p$ prime. These properties transfer to the analogous properties regarding the collection $\left(T_{k}^{\mathrm{dist}}\right)$. For clarity, we quote now (same reference) the following result, the proof of which will be reexamined in Theorem 5.1.
Proposition 4.2.
In the case when the distribution ${\mathfrak{N}}_{\chi,i\lambda}$, as defined in (4.1), is automorphic, it is automatically a Hecke distribution, by which we mean a joint eigendistribution of the collection of Hecke operators $T_{k}^{\mathrm{dist}}$. One has $T_{-1}^{\mathrm{dist}}{\mathfrak{N}}_{\chi,i\lambda}=(-1)^{\delta}\,{\mathfrak{N}}_{\chi,i\lambda}$ and, for $p$ prime,
$T_{p}^{\mathrm{dist}}\,{\mathfrak{N}}_{\chi,i\lambda}=\left[\,\chi(p)\,p^{-\frac{i\lambda}{2}}+\chi(p^{-1})\,p^{\frac{i\lambda}{2}}\right]{\mathfrak{N}}_{\chi,i\lambda}.$ (4.9)
We turn now to the definition of Eisenstein distributions, the notion in modular distribution theory to substitute for that of Eisenstein series of the non-holomorphic type. The treatment of these distributions is quite similar to that of Hecke distributions, with two extra terms added, and done in detail in [20, Section 1.1]: we may thus satisfy ourselves with recalling their basic properties.
For ${\mathrm{Re\,}}\nu<0,\,\nu\neq-1$, and $h\in{\mathcal{S}}(\mathbb{R}^{2})$, one defines ([19, p.93] or [20, p.15])
$\frac{1}{2}\,\,\big\langle{\mathfrak{E}}_{\nu}\,,\,h\big\rangle\,=\frac{1}{2}\,\zeta(-\nu)\,\int_{-\infty}^{\infty}|t|^{-\nu-1}\left({\mathcal{F}}_{1}^{-1}h\right)(0,t)\,dt\\ +\frac{1}{2}\,\zeta(1-\nu)\,\int_{-\infty}^{\infty}|t|^{-\nu}h(t,0)\,dt+\frac{1}{2}\,\sum_{n\neq 0}\sigma_{\nu}(n)\,\int_{-\infty}^{\infty}|t|^{-\nu-1}\left({\mathcal{F}}_{1}^{-1}h\right)\left(\frac{n}{t},\,t\right)\,dt,$ (4.10)
where $\sigma_{\nu}(n)=\sum_{1\leq d|\,n}d^{\nu}$. After the power function $t\mapsto|t|^{\mu}$ has been given a meaning, as a distribution on the line, for $\mu\neq-1,-3,\dots$, this decomposition is actually valid for $\nu\neq\pm 1,\,\,\nu\neq 0$. There is no need to exclude $\nu=0$ if one does not separate the first two terms on the right-hand side of (4.10). One shows that, as a tempered distribution, ${\mathfrak{E}}_{\nu}$ extends as an analytic function of $\nu$ for $\nu\neq\pm 1$, and that ${\mathrm{Res}}_{\nu=-1}{\mathfrak{E}}_{\nu}=-1,\,\,{\mathrm{Res}}_{\nu=1}{\mathfrak{E}}_{\nu}=\delta$, the unit mass at the origin of $\mathbb{R}^{2}$.
For a change, let us recall the simple proof of the following, which gives the transfer under $\Theta$ of Hecke distributions or Eisenstein distributions. We first rewrite (4.1) as
[…]
with
$\phi(k)=\sum_{mn=k}\,\chi\left(\frac{m}{n}\right)\,|n|^{i\lambda},\qquad k\in\mathbb{Z}^{\times}$ (4.12)
(note that $\phi(1)=2$).
Proposition 4.3.
One has
$\left(\Theta\,{\mathfrak{N}}_{\chi,i\lambda}\right)(x+iy)=\frac{1}{2}\,y^{\frac{1}{2}}\sum_{k\neq 0}\phi(k)\,K_{\frac{i\lambda}{2}}(2\pi\,|k|\,y)\,e^{2i\pi kx}.$ (4.13)
Similarly, setting $\zeta^{*}(s)=\pi^{-\frac{s}{2}}\Gamma(\frac{s}{2})\,\zeta(s)$, one has
$\left(\Theta\,{\mathfrak{E}}_{\nu}\right)(x+iy)=\zeta^{*}(1-\nu)\,y^{\frac{1-\nu}{2}}+\zeta^{*}(1+\nu)\,y^{\frac{1+\nu}{2}}\\ +2\,y^{\frac{1}{2}}\sum_{k\neq 0}|k|^{-\frac{\nu}{2}}\,\sigma_{\nu}(|k|)\,K_{\frac{\nu}{2}}(2\pi\,|k|\,y)\,e^{2i\pi kx}.$ (4.14)
One has ${\mathcal{F}}^{\mathrm{symp}}\,{\mathfrak{N}}_{\chi,i\lambda}={\mathfrak{N}}_{\chi^{-1},-i\lambda}$ and ${\mathcal{F}}^{\mathrm{symp}}{\mathfrak{E}}_{\nu}={\mathfrak{E}}_{-\nu}$.
Proof.
Making the Fourier expansions (1.2) and (4.10) more explicit, as
${\mathfrak{N}}_{\chi,i\lambda}(x,\xi)$ $=\frac{1}{4}\sum_{k\neq 0}\phi(k)\,|\xi|^{-1-i\lambda}\exp\left(\frac{2i\pi\,kx}{\xi}\right),$
${\mathfrak{E}}_{\nu}(x,\xi)$ $=\zeta(-\nu)\,|\xi|^{-\nu-1}+\zeta(1-\nu)\,|x|^{-\nu}\delta(\xi)+\sum_{k\neq 0}\sigma_{\nu}(k)\,|\xi|^{-\nu-1}\exp\left(\frac{2i\pi\,kx}{\xi}\right),$ (4.15)
and it suffices to compute the Theta-transform (3.1) of the function $h_{\nu,k}(x,\xi)=|\xi|^{-\nu-1}\exp\left(\frac{2i\pi\,kx}{\xi}\right)$. Setting $z=\alpha+iy$ (the use of $x$ has been preempted), one first computes the (Gaussian) $dx$-integral, obtaining
$(\Theta\,h_{\nu,k})(\alpha+iy)=e^{2i\pi k\alpha}y^{\frac{1}{2}}\int_{-\infty}^{\infty}|\xi|^{-\nu-1}\exp\left(-\pi\,y(\xi^{2}+\xi^{-2})\right)\,d\xi,$ (4.16)
after which, using a standard integral definition [12, p.85] of modified Bessel functions, one obtains
$(\Theta\,h_{\nu,k})(x+iy)=2\,y^{\frac{1}{2}}K_{\frac{\nu}{2}}(2\pi\,|k|\,y)\,e^{2i\pi kx}.$ (4.17)
This computation was done in [20, p.116], but a serious reason to change the normalization (by a “factor” $2^{\frac{1}{2}+i\pi{\mathcal{E}}}$) when dealing with arithmetic has been explained in [21, Section 6.1]. There is no difficulty in summing the $k$-series, thanks to the exponential decrease of the function $K_{\frac{\nu}{2}}$.
The last assertion follows from the identity ${\mathcal{F}}^{\mathrm{symp}}h_{\nu,k}=|k|^{-\nu}h_{-\nu,k}$, the verification of which is straightforward.
The function $\Theta\,{\mathfrak{N}}_{\chi,i\lambda}$ is automorphic, hence a cusp-form in view of its expansion (4.13): from (4.9), it is a Hecke eigenform, normalized in Hecke’s way (a notion recalled in (4.18) below) since $\phi(1)=2$. The right-hand side of (4.14) is the familiar-looking Fourier expansion [6, p.66] of the function $\zeta^{*}(\nu)\,E_{\frac{1-\nu}{2}}(z)$, where $E_{\frac{1-\nu}{2}}$ is the non-holomorphic Eisenstein series so denoted. To complete the picture, we must show that every Hecke-normalized Hecke eigenform is of the form $\Theta\,{\mathfrak{N}}_{\chi,i\lambda}$ for some choice of the pair $(\chi,i\lambda)$. This was done in [20, Prop.2.1.1] and [20, Theor.2.1.2].
Proposition 4.4.
Let $\mathcal{N}$} be a Hecke eigenform, normalized in Hecke’s way, with Fourier expansion
${\mathcal{N}}(x+iy)=y^{\frac{1}{2}}\sum_{k\neq 0}b_{k}\,K_{\frac{i\lambda}{2}}(2\pi\,|k|\,y)\,e^{2i\pi kx},$ (4.18)
normalized in Hecke’s way (i.e., $b_{1}=1$): define $\delta=0$ or $1$ according to the parity of ${\mathcal{N}}(z)$ under the map $z\mapsto-\overline{z}$. Choose for $\lambda$ any of the two square roots of $\lambda^{2}$ ($K_{\frac{i\lambda}{2}}$ depends only $\lambda^{2}$). Then, choosing for every prime $p$ any of the two solutions of the equation $b_{p}=\theta_{p}+\theta_{p}^{-1}$, define $\chi$ as the unique character of $\mathbb{Q}^{\times}$ such that $\chi(p)=p^{\frac{i\lambda}{2}}\theta_{p}$ for $p$ prime and $\chi(-1)=(-1)^{\delta}$. The distribution
${\mathfrak{N}}(x,\xi)=\frac{1}{2}\sum_{k\neq 0}b_{k}\,\,|k|^{\frac{i\lambda}{2}}\,\,|\xi|^{-1-i\lambda}\exp\left(\frac{2i\pi kx}{\xi}\right)$ (4.19)
coincides with the distribution ${\mathfrak{N}}_{\chi,i\lambda}$, as defined in (4.1). It is a Hecke distribution, and one has
$\Theta\,{\mathfrak{N}}_{\chi,i\lambda}={\mathcal{N}}.\\$ (4.20)
One point was, however, missing in the given reference. Setting
$\Lambda(s,\,{\mathcal{N}})=\pi^{-s}\Gamma\left(\frac{s+\delta}{2}+\frac{i\lambda}{4}\right)\Gamma\left(\frac{s+\delta}{2}-\frac{i\lambda}{4}\right)L(s,\,{\mathcal{N}})$ (4.21)
with
$L(s,\,{\mathcal{N}})=\sum_{k\geq 1}b_{k}k^{-s}=\prod_{p}\left(1-b_{p}p^{-s}+p^{-2s}\right)^{-1},$ (4.22)
the function $\Lambda(s,\,{\mathcal{N}})$ extends as an entire function of $s$, satisfying the functional equation [1, p.107]
$\Lambda(s,\,{\mathcal{N}})=(-1)^{\delta}\Lambda(1-s,\,{\mathcal{N}})$ (4.23)
(in the reference just given, the parameter $\nu$ is the one denoted here as $\frac{i\lambda}{2}$). The function $L^{\natural}\left(s,\,{\mathfrak{N}}_{\chi,i\lambda}\right)$ does not coincide with $\Lambda(s,\,{\mathcal{N}})$: of course, it has to be more precise, since it distinguishes between $\lambda$ and $\lambda$. But the two are linked by the identity
$L^{\natural}\left(s,\,{\mathfrak{N}}_{\chi,i\lambda}\right)=\frac{1}{2}\,\frac{(-i)^{\delta}\pi^{\frac{1-i\lambda}{2}}}{\Gamma\left(\frac{s+\delta}{2}-\frac{i\lambda}{4}\right)\Gamma\left(\frac{1-s+\delta}{2}-\frac{i\lambda}{4}\right)}\,\Lambda(s,\,{\mathcal{N}}).$ (4.24)
This proves that the function $L^{\natural}\left(\,\centerdot\,,\,{\mathfrak{N}}_{\chi,i\lambda}\right)$ satisfies the same functional equation as the function $\Lambda(\,\centerdot\,,\,{\mathcal{N}})$. We forgot in the given reference to show that the function $L^{\natural}\left(s,\,{\mathfrak{N}}_{\chi,i\lambda}\right)$ is polynomially bounded in vertical strips, which does not follow from the analogous fact relative to the function $\Lambda(s,\,{\mathcal{N}})$. The proof of [1, Lemma 1.9.1] applies with the following modification. Replace the integral representation of the function $K_{\frac{i\lambda}{2}}$ used in the given reference by the pair (assuming that $k>0$ and, just as in [1], that $y\geq 1$), to be found in [12, p.85],
$K_{\frac{i\lambda}{2}}(2\pi ky)=\int_{0}^{\infty}e^{-2\pi ky\cosh t}\cos\frac{\lambda t}{2}\,dt\\ =\pi^{-\frac{1}{2}}\Gamma(\frac{1+i\lambda}{2})(\pi ky)^{-\frac{i\lambda}{2}}\int_{0}^{\infty}(\cosh t)^{-i\lambda}\cos(\pi ky\sinh t)\,dt.$ (4.25)
The first equation yields $|K_{\frac{i\lambda}{2}}(2\pi ky)|\leq C\,e^{-\pi ky}$ and, transforming the integral in the second equation to
$\int_{0}^{\infty}\frac{d}{dt}\left[\frac{1}{\pi ky}\,(\cosh t)^{-1-i\lambda}\right]\sin(\pi ky\sinh t)\,dt,$ (4.26)
one obtains
$|K_{\frac{i\lambda}{2}}(2\pi ky)|\leq\pi^{-\frac{4}{2}}\sqrt{1+\lambda^{2}}\,\left|\Gamma\left(\frac{1+i\lambda}{2}\right)\right|\,(\pi ky)^{-1}.$ (4.27)
One makes about the best of the two estimates if, when dealing with the $k$-series (4.18) that defines ${\mathcal{N}}(iy)$, one uses the first estimate when $ky>\frac{|\lambda|}{4}$, the second when $ky<\frac{|\lambda|}{4}$ : (2.4) implies the desired result.
Recall [7, p.372] that, with $\mathcal{N}$} given by (4.18) and $\mathfrak{N}$} by (4.19), one has for $p$ prime
$T_{p}{\mathcal{N}}=b_{p}\,{\mathcal{N}}\qquad{\mathrm{and}}\qquad T_{p}^{\mathrm{dist}}{\mathfrak{N}}=b_{p}\,{\mathfrak{N}}.$ (4.28)
The equation (4.9) proves the equivalence between the set of conditions $|b_{p}|\leq 2$ and the unitarity of the character $\chi$.
5. Hecke operators and their powers
Theorem 5.1.
Given a prime $p$, and a modular distribution ${\mathfrak{N}}={\mathfrak{E}}_{\nu}$ or ${\mathfrak{N}}_{\chi,i\lambda}$, one has the identity
$\left(T_{p}^{\mathrm{dist}}\,{\mathfrak{N}}\right)(x,\xi)=\left(p^{-\frac{1}{2}+i\pi{\mathcal{E}}^{\natural}}\,{\mathfrak{N}}\right)(x,\xi)+\frac{1}{p}\sum_{b\,{\mathrm{mod}}\,p}\left(p^{\frac{1}{2}-i\pi{\mathcal{E}}^{\natural}}\,{\mathfrak{N}}\right)(x+b\xi,\,\xi).$ (5.1)
Proof.
From (4), one has, defining $\widetilde{\phi}(\lambda)=\phi(\lambda)$ if $\lambda\in\mathbb{Z}^{\times}$ (cf. (4.12)), $\widetilde{\phi}(\lambda)=0$ if $\lambda\in\mathbb{Q}$ is zero or fails to be an integer,
[…]
If, in the second formula, we substitute $x+b\xi$ for $x$ and perform $\frac{1}{p}$ times the summation with respect to $b$, we come across the sum $\frac{1}{p}\sum_{b\,{\mathrm{mod}}\,p}\exp\left(\frac{2i\pi kb}{p}\right)={\mathrm{char}}(k\equiv 0\,{\mathrm{mod}}\,p)$. It follows that the right-hand side of (5.1) is
[…]
For $k\in\mathbb{Z}^{\times}$, one has (in what follows, $(m,n)$ denotes the pair of integers $m,n$, not their g.c.d. !))
$\{(m,n)\colon mn=pk\}=\{\,(pm_{1},n)\colon m_{1}n=k\}\,\cup\,\{\,(m,pn_{1})\colon mn_{1}=k\},$ (5.4)
not a disjoint union if $p|k$: in such a case, the two sets intersect along the set $\{\,(pm_{1},\,pn_{1})\colon m_{1}n_{1}=\frac{k}{p}\}$, which leads to the equation
$\phi(pk)=\left[\chi(p)+p^{i\lambda}\chi(p^{-1})\right]\phi(k)-p^{i\lambda}\,\widetilde{\phi}\left(\frac{k}{p}\right),\qquad k\in\mathbb{Z}^{\times}.$ (5.5)
It follows that
$p^{-\frac{i\lambda}{2}}\phi(pk)+p^{\frac{i\lambda}{2}}\widetilde{\phi}\left(\frac{k}{p}\right)=\left[p^{-\frac{i\lambda}{2}}\chi(p)+p^{\frac{i\lambda}{2}}\chi(p^{-1})\right]\phi(k),$ (5.6)
and the right-hand side of (5.1) agrees with
$\left(p^{\frac{i\lambda}{2}}(\chi(p))^{-1}+p^{-\frac{i\lambda}{2}}\chi(p)\right)\,{\mathfrak{N}}_{\chi,i\lambda}(x,\xi),$ (5.7)
the same as $\left(T_{p}^{\mathrm{dist}}\,{\mathfrak{N}}_{\chi,i\lambda}\right)(x,\xi)$ in view of (4.9).
In the case of the Eisenstein distribution, just replacing $\chi$ by $\chi_{\overset{}{0}}$ everywhere would seem to lead to the fact that nothing is changed in the proof. The result is correct, but one must not, in this case, forget the two extra terms in (4.10), which are multiples of $|\xi|^{-\nu-1}$ and $|x|^{-\nu}\delta(\xi)$, Now, one has
$p^{-\frac{1}{2}+i\pi{\mathcal{E}}^{\natural}}\left(|\xi|^{-\nu-1}\right)$ $=p^{\frac{\nu}{2}}\,|\xi|^{-\nu-1},\qquad$ $p^{-\frac{1}{2}+i\pi{\mathcal{E}}^{\natural}}\left(|x|^{-\nu}\delta(\xi)\right)$ $=p^{-\frac{\nu}{2}}\,|x|^{-\nu}\delta(\xi),$
$p^{\frac{1}{2}-i\pi{\mathcal{E}}^{\natural}}\left(|\xi|^{-\nu-1}\right)$ $=p^{-\frac{\nu}{2}}\,|\xi|^{-\nu-1},\qquad$ $p^{\frac{1}{2}-i\pi{\mathcal{E}}^{\natural}}\left(|x|^{-\nu}\delta(\xi)\right)$ $=p^{\frac{\nu}{2}}\,|x|^{-\nu}\delta(\xi),$ (5.8)
and one obtains the desired result, noting that on the special terms under consideration here, the translation $x\mapsto x-b\xi$ has no effect.
Simplify $T_{p}^{\mathrm{dist}}$ as $T$ and set $R=p^{-\frac{1}{2}+i\pi{\mathcal{E}}^{\natural}}$. The operators $R$ and $R^{-1}$ act on arbitrary (tempered) distributions. Given $j\in\mathbb{Z}$, denote as $\Gamma_{\infty}^{\left(p^{j}\right)}$ the subgroup of $SL(2,\mathbb{R})$ generated by the matrix $\left(\begin{smallmatrix}1&p^{j}\\ 0&1\end{smallmatrix}\right)$, and denote as ${\mathrm{Inv}}(p^{j})$ the linear space of tempered distributions $\mathfrak{S}$} invariant under the action of $\Gamma_{\infty}^{\left(p^{j}\right)}$, i.e., satisfying the “periodicity” condition ${\mathfrak{S}}(x+p^{j}\xi,\xi)={\mathfrak{S}}(x,\xi)$. We shall express by the notation $Q_{1}\sim Q_{2}$ the fact that the operators $Q_{1}$ and $Q_{2}$ are well-defined on ${\mathrm{Inv}}(1)$ (or on larger spaces) and agree there.
Lemma 5.2.
For $j,\ell\in\mathbb{Z}$, the operator $R^{\ell}$ maps ${\mathrm{Inv}}(p^{j})$ into ${\mathrm{Inv}}(p^{j-\ell})$: in particular, if $\ell\geq 0$, it preserves the space ${\mathrm{Inv}}(1)$. Setting, for $\gamma\in\mathbb{R}$, $\tau[\gamma]=\exp\left(\gamma\,\xi\frac{\partial}{\partial x}\right)$, so that $\left(\tau[\gamma]\,{\mathfrak{S}}\right)(x,\xi)={\mathfrak{S}}(x+\gamma\xi,\,\xi)$, define when $r$ and $\ell$ are non-negative integers the operators
$\sigma_{r}^{(\ell)}=p^{-r}\sum_{b\,{\mathrm{mod}}\,p^{r}}\tau\left[b\,p^{\ell-r}\right],\qquad\sigma_{r}=\sigma_{r}^{0}.$ (5.9)
The first operator sends the space ${\mathrm{Inv}}(p^{\ell})$ into ${\mathrm{Inv}}(p^{\ell-r})$ : in particular, $\sigma_{r}$ sends ${\mathrm{Inv}}(1)$ to ${\mathrm{Inv}}(p^{-r})$. For $r\geq 0,\,\ell\geq 0$, one has
$R^{\ell}\sigma_{r}\sim\sigma_{\ell+r}\,R^{\ell},\qquad R^{-\ell}\sigma_{r}\sim\sigma_{r}^{(\ell)}\,R^{-\ell}.$ (5.10)
If $r\geq 0$, one has $\sigma_{r}\,R^{-1}\sigma_{1}\sim R^{-1}\sigma_{r+1}$. The operator $\sigma_{r}^{(\ell)}$ commutes with ${\mathcal{F}}^{\mathrm{symp}}$ for all values of $r,\ell$.
Proof.
One has for every distribution $\mathfrak{S}$}
$(R^{\ell}{\mathfrak{S}})\left(x+p^{j-\ell},\,\xi\right)$ $=p^{-\frac{\ell}{2}}{\mathfrak{S}}\left(p^{\frac{\ell}{2}}\left(x+p^{j-\ell}\xi\right),\,p^{-\frac{\ell}{2}}\xi\right)$
$=p^{-\frac{\ell}{2}}{\mathfrak{S}}\left(p^{\frac{\ell}{2}}x+p^{j}\left(p^{-\frac{\ell}{2}}\xi\right),\,p^{-\frac{\ell}{2}}\xi\right),$ (5.11)
so that $R^{\ell}{\mathfrak{S}}$ is invariant under $\Gamma_{\infty}^{(p^{j-\ell})}$ if $\mathfrak{S}$} invariant under $\Gamma_{\infty}^{(p^{j})}$.
It is immediate that, for every $\gamma\in\mathbb{R}$, one has
$R\,\tau[\gamma]=\tau\left[\frac{\gamma}{p}\right]\,R,\qquad R^{-1}\tau[\gamma]=\tau[p\gamma]\,R^{-1}.$ (5.12)
One has
$(R^{\ell}\sigma_{r}\,{\mathfrak{S}})(x,\xi)$ $=R^{\ell}\,\left[(x,\xi)\mapsto\,p^{-r}\sum_{b\,{\mathrm{mod}}\,p^{r}}{\mathfrak{S}}\left(x+\frac{b\,\xi}{p^{r}},\,\xi\right)\,\right]$
$=p^{-r-\frac{\ell}{2}}\sum_{b\,{\mathrm{mod}}\,p^{r}}{\mathfrak{S}}\left(p^{\frac{\ell}{2}}x+b\,p^{-\frac{\ell}{2}-r}\xi,\,p^{-\frac{\ell}{2}}\xi\right)$ (5.13)
and
$(\sigma_{\ell+r}R^{\ell}\,{\mathfrak{S}})(x,\xi)$ $=\sigma_{\ell+r}\left[(x,\xi)\mapsto p^{-\frac{\ell}{2}}{\mathfrak{S}}\left(p^{\frac{\ell}{2}}x,\,p^{-\frac{\ell}{2}}\xi\right)\right]$
$=p^{-r-\frac{3\ell}{2}}\sum_{b\,{\mathrm{mod}}\,p^{\ell+r}}{\mathfrak{S}}\,\left(p^{\frac{\ell}{2}}\left(x+\frac{b\,\xi}{p^{\ell+r}}\right),\,p^{-\frac{\ell}{2}}\xi\right).$ (5.14)
In the two formulas, the arguments are the same, but the domains of averaging are not. However, when two classes $b$ mod $p^{\ell+r}$ agree mod $p^{r}$, the first arguments of the two expressions differ by a multiple of $p^{-\frac{\ell}{2}}\xi$, so that the two expressions are identical if ${\mathfrak{S}}\in{\mathrm{Inv}}(1)$.
Next,
$(R^{-\ell}\sigma_{r}\,{\mathfrak{S}})(x,\xi)$ $=R^{-\ell}\left[(x,\xi)\mapsto p^{-r}\sum_{b\,{\mathrm{mod}}\,p^{r}}{\mathfrak{S}}\left(x+\frac{b\xi}{p^{r}},\,\xi\right)\right]$
$=p^{\frac{\ell}{2}-r}\sum_{b\,{\mathrm{mod}}\,p^{r}}{\mathfrak{S}}\left(p^{-\frac{\ell}{2}}x+p^{\frac{\ell}{2}-r}\xi,\,p^{\frac{\ell}{2}}\xi\right)$
$=p^{-r}\sum_{b\,{\mathrm{mod}}\,p^{r}}(R^{-\ell}{\mathfrak{S}})\left(x+b\,p^{\ell-r}\xi,\,\xi\right).$ (5.15)
This is the same as $\left(\sigma_{r}^{(\ell)}\,R^{-\ell}{\mathfrak{S}}\right)(x,\xi)$.
Finally, if ${\mathfrak{S}}\in{\mathrm{Inv}}(1)$, ${\mathfrak{T}}=\sigma_{1}{\mathfrak{S}}\in{\mathrm{Inv}}(p^{-1})$, and one observes first that both operators $\sigma_{r}\,R^{-1}\sigma_{1}$ and $R^{-1}\sigma_{r+1}$ are well-defined on ${\mathrm{Inv}}(1)$. One has
$(\sigma_{r}\,R^{-1}{\mathfrak{T}})(x,\xi)$ $=\sigma_{r}\left[(x,\xi)\mapsto p^{\frac{1}{2}}\,{\mathfrak{T}}\left(p^{-\frac{1}{2}}x,\,p^{\frac{1}{2}}\xi\right)\right]$
$=p^{-r+\frac{1}{2}}\sum_{b\,{\mathrm{mod}}\,p^{r}}{\mathfrak{T}}\left(p^{-\frac{1}{2}}x+b\,p^{-r-\frac{1}{2}}\xi,\,p^{\frac{1}{2}}\xi\right)$
$=R^{-1}\left[p^{-r}\sum_{b\,{\mathrm{mod}}\,p^{r}}{\mathfrak{T}}\left(x+\frac{b\xi}{p^{r+1}},\,\xi\right)\right].$ (5.16)
Then,
$(\sigma_{r}\,R^{-1}\sigma_{1}\,{\mathfrak{S}})(x,\xi)=R^{-1}\bigg[p^{-r-1}\sum_{\begin{array}[]{c}b\,{\mathrm{mod}}\,p^{r}\\ \beta\,{\mathrm{mod}}\,p\end{array}}{\mathfrak{S}}\left(x+\frac{b\xi}{p^{r+1}}+\frac{\beta\xi}{p},\,\xi\right)\bigg].$ (5.17)
As $b$ and $\beta$ run through the classes indicated as a subscript, $b+p^{r}\beta$ describes a full set of classes modulo $p^{r+1}$, so that the right-hand side is indeed $(R^{-1}\sigma_{r+1}{\mathfrak{S}})(x,\xi)$.
The last assertion follows from the fact that ${\mathcal{F}}^{\mathrm{symp}}{\mathfrak{S}}\in{\mathrm{Inv}}(p^{\ell})$ if ${\mathfrak{S}}\in{\mathrm{Inv}}(p^{\ell})$ and from the identity
$\sigma_{r}^{(\ell)}=p^{-r}\sum_{b\,{\mathrm{mod}}\,p^{r}}\exp\left(b\,p^{\ell-r}\xi\frac{\partial}{\partial x}\right),$ (5.18)
if one notes that the operator $\xi\frac{\partial}{\partial x}$ commutes with ${\mathcal{F}}^{\mathrm{symp}}$.
In (5.1), the operator $T=T_{p}^{\mathrm{dist}}$ has been obtained, as an operator of modular distributions, as
$T=R+\sigma_{1}^{(1)}R^{-1}=R+R^{-1}\sigma_{1}$ (5.19)
(recall that $R=p^{-\frac{1}{2}+i\pi{\mathcal{E}}^{\natural}}$). The first equation can be verified from the definition of $\sigma_{r}^{(\ell)}$, and the second then follows from Lemma 5.2. This equation is still a valid definition when applied to distributions in the space ${\mathrm{Inv}}(1)$, and we shall use its application to automorphic distributions in the last two sections.
Proposition 5.3.
Given $k=1,2,\dots$ and $\ell$ such that $0\leq\ell\leq k$, there are non-negative integers $\alpha_{k,\ell}^{(0)},\,\alpha_{k,\ell}^{(1)},\,\dots,\,\alpha_{k,\ell}^{(\ell)}$ , satisfying the conditions:
(i) $\alpha_{k,\ell}^{(0)}+\alpha_{k,\ell}^{(1)}+\dots+\alpha_{k,\ell}^{(\ell)}=\left(\begin{smallmatrix}k\\ \ell\end{smallmatrix}\right)$ for all $k,\ell$ ,
(ii) $2\ell-k-r\leq 0$ whenever $\alpha_{k,\ell}^{(r)}\neq 0$,
such that one has the identity (between two operators on ${\mathrm{Inv}}(1)$)
$T^{k}=\sum_{\ell=0}^{k}R^{k-2\ell}\,\left(\alpha_{k,\ell}^{(0)}\,I+\alpha_{k,\ell}^{(1)}\,\sigma_{1}+\dots+\alpha_{k,\ell}^{(\ell)}\,\sigma_{\ell}\right).$ (5.20)
Proof.
By induction. Assuming that the given formula holds, we write $T^{k+1}=T^{k}(R+R^{-1}\sigma_{1})$ and we use the equations $\sigma_{r}\,R\sim R\,\sigma_{r-1}$ ($r\geq 1$) and $\sigma_{r}\,R^{-1}\sigma_{1}\sim R^{-1}\sigma_{r+1}$ at the end of Lemma 5.2. We obtain
$T^{k+1}$ $=\sum_{\ell=0}^{k}R^{k+1-2\ell}\left(\alpha_{k,\ell}^{(0)}\,I+\alpha_{k,\ell}^{(1)}\,I+\alpha_{k,\ell}^{(2)}\,\sigma_{1}+\dots+\alpha_{k,\ell}^{(\ell)}\,\sigma_{\ell-1}\right)$
$+\sum_{\ell=0}^{k}R^{k-1-2\ell}\left(\alpha_{k,\ell}^{(0)}\,\sigma_{1}+\alpha_{k,\ell}^{(1)}\,\sigma_{2}+\dots+\alpha_{k,\ell}^{(\ell)}\,\sigma_{\ell+1}\right),$ (5.21)
or
$T^{k+1}$ $=\sum_{\ell=0}^{k}R^{k+1-2\ell}\left(\alpha_{k,\ell}^{(0)}\,I+\alpha_{k,\ell}^{(1)}\,I+\alpha_{k,\ell}^{(2)}\,\sigma_{1}+\dots+\alpha_{k,\ell}^{(\ell)}\,\sigma_{\ell-1}\right)$
$+\sum_{\ell=1}^{k+1}R^{k+1-2\ell}\left(\alpha_{k,\ell-1}^{(0)}\,\sigma_{1}+\alpha_{k,\ell-1}^{(1)}\,\sigma_{2}+\dots+\alpha_{k,\ell-1}^{(\ell-1)}\,\sigma_{\ell}\right).$ (5.22)
The point (i) follows, using $\left(\begin{smallmatrix}k\\ \ell\end{smallmatrix}\right)+\left(\begin{smallmatrix}k\\ \ell-1\end{smallmatrix}\right)=\left(\begin{smallmatrix}k+1\\ \ell\end{smallmatrix}\right)$. Next, looking again at (5), one observes that, in the expansion of $T^{k+1}$, the term $R^{k+1-2\ell}\sigma_{r}$ is the sum of two terms originating (in the process of obtaining $T^{k+1}$ from $T^{k}$) from the terms $R^{k-2\ell}\sigma_{r+1}$ and $R^{k-2\ell+2}\sigma_{r-1}$. The condition $2\ell-(k+1)-r\leq 0$ is certainly true if either $2\ell-k-(r+1)\leq 0$ or $(2\ell-2)-k-(r-1)\leq 0$, which proves the point (ii) by induction. Since $\sigma_{r}{\mathfrak{S}}\in{\mathrm{Inv}}\left(p^{-r}\right)$ if ${\mathfrak{S}}\in{\mathrm{Inv}}(1)$, this condition implies that $R^{k-2\ell}\sigma_{r}$ preserves the space ${\mathrm{Inv}}(1)$.
Remark 5.1. As mentioned in the introduction, using the $N$th power of the Hecke operator is not truly necessary for the aim of proving the Ramanujan-Petersson conjecture, and one could thus satisfy oneself with the case $N=1$. If so doing, one might dispense with Proposition 5.3, but most of Lemma 5.2 would still be needed.
6. A generating object for modular distributions
A basic collection of distributions is made of the “elementary” line measures
${\mathfrak{s}}^{a}_{b}(x,\xi)=e^{2i\pi ax}\delta(\xi-b),\qquad a,b\in\mathbb{R}.$ (6.1)
Superpositions of these conduct in a natural way to modular distributions of interest. For instance, with $\chi$ as in Section 4, consider the series
${\mathfrak{T}}_{\chi}=\pi\sum_{m,n\neq 0}\chi\left(\frac{m}{n}\right)\,{\mathfrak{s}}^{m}_{n}.$ (6.2)
The distribution ${\mathfrak{T}}_{\chi}$ decomposes [21, p.116] into homogeneous components as $\int_{-\infty}^{\infty}{\mathfrak{N}}_{\chi,i\lambda}\,d\lambda$, with ${\mathfrak{N}}_{\chi,i\lambda}$ as defined in (4.1). If one replaces $\chi$ by the trivial character $\chi_{0}=1$, and one replaces the domain of summation by the one defined by the condition $|m|+|n|\neq 0$, the same decomposition will lead instead to the Eisenstein distributions ${\mathfrak{E}}_{i\lambda}$.
For $j=0,1,\dots$, define the distribution
$\left({\mathfrak{s}}_{1}^{1}\right)^{j}=\pi^{2}{\mathcal{E}}^{2}(\pi^{2}{\mathcal{E}}^{2}+1)\dots(\pi^{2}{\mathcal{E}}^{2}+(j-1)^{2})\,{\mathfrak{s}}_{1}^{1}.$ (6.3)
It was shown in [18, Theorem 3.3] that for $j\geq 1$, setting $\Gamma_{\infty}=\{\left(\begin{smallmatrix}1&b\\ 0&1\end{smallmatrix}\right)\colon b\in\mathbb{Z}\}$, the series
${\mathfrak{B}}^{j}=\frac{1}{2}\sum_{g\in\Gamma/\Gamma_{\infty}}\left({\mathfrak{s}}_{1}^{1}\right)^{j}\,\circ\,g^{-1}$ (6.4)
converges in the space ${\mathcal{S}}^{\prime}(\mathbb{R}^{2})$ provided that, before summation, one always groups the terms corresponding to $g=\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right)$ or $g=\left(\begin{smallmatrix}-a&-b\\ -c&-d\end{smallmatrix}\right)$. As such, it is an even distribution (i.e., giving zero when tested on globally odd functions), characterized by its restriction as a continuous linear form on ${\mathcal{S}}_{\mathrm{even}}(\mathbb{R}^{2})$.
The distribution ${\mathfrak{B}}^{j}$ is invariant under ${\mathcal{F}}^{\mathrm{symp}}$, and it follows (3.4) that it carries exactly the same information as its image $f^{j}$ under $\Theta$ (cf.(6.6) below), a well-known automorphic function introduced by Selberg [14]. However, it is the automorphic distribution version that will make possible the analysis in Sections 7 and 8 (the hard part of the paper), to begin with integrations by parts. On the other hand, we shall, as a start, just transfer known results of automorphic function theory to obtain the spectral decomposition of ${\mathfrak{B}}^{j}$.
For the benefit of readers not familiar with automorphic function theory, and the necessities of notation, here is a quite short summary of classical results. Nice presentations of this theory (accessible to non-experts, including the present author) are to be found in [1, 6, 7] and elsewhere.
The space $\)L^{2}(\Gamma\backslash\mathbb{H}$) is built with the help of the $SL(2,\mathbb{R})$-invariant measure $dm(z)=y^{-2}dx\,dy$ (with $z=x+iy$) on $\mathbb{H}$ and of a fundamental domain $D$ for the action of $\Gamma$ in $\mathbb{H}$, say $D=\{z\colon|z|>1,\,|{\mathrm{Re\,}}z|<\frac{1}{2}\}$. A suitable self-adjoint realization of $\Delta$, in this space, exists, and all Hecke operators $T_{k}$, too, are self-adjoint there. The Eisenstein series $E_{\frac{1-i\lambda}{2}}$ make up a complete set of generalized eigenfunctions for the continuous part of the spectrum of $\Delta$, while the Hecke eigenforms are genuine eigenvectors (they lie in $L^{2}(\Gamma\backslash\mathbb{H})=L^{2}(D)$) of this operator. Both families are by definition made of joint eigenfunctions of all Hecke operators. Letting after some work the theory of compact self-adjoint operators play, one finds that the true eigenvalues make up a sequence $\left(\frac{1+\lambda_{r}^{2}}{4}\right)_{r\geq 1}$ going to infinity: that there are indeed infinitely many eigenvalues is subtler and relies on Selberg’s trace formula. For every $r\geq 1$, the space of Hecke eigenforms corresponding to this eigenvalue is finite-dimensional, generated by some family $\left({\mathcal{N}}^{r,\iota}\right)_{1\leq\iota\leq d_{r}}$ of pairwise orthogonal Hecke eigenforms.
The Hecke normalization of Hecke eigenforms ${\mathcal{N}}^{r,\iota}$, for which $b_{1}=1$ (cf. Proposition 4.4), has the property that the values the Hecke operators take on ${\mathcal{N}}^{r,\iota}$ can be read directly on the Fourier coefficients of this Hecke eigenform [6, p.128]. But since the normalization of ${\mathcal{N}}^{r,\iota}$ has already been chosen, it cannot be expected (and it is considerably far from being the case [15]) that it could be normalized in the space $\)L^{2}(\Gamma\backslash\mathbb{H}$) as well: we shall thus introduce the norm there, denoted as $\|{\mathcal{N}}^{r,\iota}\|$, of ${\mathcal{N}}^{r,\iota}$. In terms of modular form theory, the Ramanujan-Petersson conjecture is the assertion that, with $b_{p}$ as defined in (4.18), one has $|b_{p}|\leq 2$ for every Hecke-normalized eigenform and every prime $p$.
[…]
Theorem 6.1.
For $j=1,2,\dots$, the automorphic distribution ${\mathfrak{B}}^{j}$ introduced in (6.4) admits the decomposition, convergent in the space of continuous linear forms on ${\mathcal{S}}_{\mathrm{even}}(\mathbb{R}^{2})$,
[…]
Proof.
It was given in [18, theor. 4.3]. Let us recall the main points. First, one proves [18, theor. 3.3] the convergence, in the sense indicated, of the series (6.3), (6.4) defining the left-hand side of (6.5), and one obtains [18, theor. 3.5] the identity
$f^{j}(z)\colon=\left(\Theta\,{\mathfrak{B}}^{j}\right)(z)=(4\pi)^{j}\frac{\Gamma(\frac{1}{2}+j)}{\Gamma(\frac{1}{2})}\\ \times\,\frac{1}{2}\sum_{\left(\begin{smallmatrix}n&n_{1}\\ m&m_{1}\end{smallmatrix}\right)\in\Gamma/\Gamma_{\infty}}\left(\frac{{\mathrm{Im\,}}z}{|-mz+n|^{2}}\right)^{j+\frac{1}{2}}\exp\left(2i\pi\frac{m_{1}z-n_{1}}{-mz+n}\right).$ (6.6)
The spectral decomposition of this automorphic function can be found in several places [3, 4]. It is
$f^{j}(z)$ $=\frac{1}{4\pi}\int_{-\infty}^{\infty}\frac{\Gamma(j-\frac{i\lambda}{2})\Gamma(j+\frac{i\lambda}{2})}{\zeta^{*}(i\lambda)\zeta^{*}(-i\lambda)}\,E^{*}_{\frac{1-i\lambda}{2}}(z)\,d\lambda$
[…]
Then, one notes that the two terms of the decomposition (6.5) (the integral and the series) are convergent in ${\mathcal{S}}^{\prime}(\mathbb{R}^{2})$, i.e., the integral and series obtained when testing it against any function in ${\mathcal{S}}(\mathbb{R}^{2})$ are absolutely convergent. Indeed, treating the integral part of the decomposition does not require more than a pair of integrations by parts. The series requires using the bound ${\mathrm{O}}(r^{\frac{1}{2}})$ for the number of $\lambda_{j}$’s with $|\lambda_{j}|<r$, a consequence of the Selberg equivalent [6, p.174] originating from the trace formula, and the estimate [15]
$\|{\mathcal{N}}^{r,\iota}\|^{-1}\leq C\,\big|\Gamma\left(\frac{i\mu_{r}}{2}\right)\big|^{-1}.$ (6.8)
Using the fact (Proposition 4.3) that (with our present notation)
$\Theta\,{\mathfrak{E}}_{\nu}=E_{\frac{1-\nu}{2}}^{*},\qquad\Theta\,{\mathfrak{N}}^{r,\iota}={\mathcal{N}}^{|r|,\iota},$ (6.9)
one obtains, since the map $\Theta$ is one-to-one when restricted the the space of ${\mathcal{F}}^{\mathrm{symp}}$-invariant distributions, the identity (6.5).
7. The operator with $(2i\pi{\mathcal{E}})^{2}\,{\mathfrak{B}}$ for symbol
We start with a review of the main representation-theoretic aspects of pseudodifferential analysis. If ${\mathfrak{S}}\in{\mathcal{S}}^{\prime}(\mathbb{R}^{2})$, the operator $\Psi({\mathfrak{S}})$ with symbol $\mathfrak{S}$} is the linear operator from ${\mathcal{S}}(\mathbb{R})$ to ${\mathcal{S}}^{\prime}(\mathbb{R})$ weakly defined by the equation
$\left(\Psi({\mathfrak{S}})\,u\right)(x)=\frac{1}{2}\,\int_{\mathbb{R}^{2}}{\mathfrak{S}}(\frac{x+y}{2},\,\xi)\,e^{i\pi(x-y)\xi}\,dy\,d\xi,$ (7.1)
in other words the operator with integral kernel
$K(x,\,y)=({\mathcal{F}}_{2}^{-1}{\mathfrak{S}})\left(\frac{x+y}{2},\,\frac{x-y}{2}\right).$ (7.2)
Yes, this definition coincides with the one given in (3.5), (3.6) in terms of the Wigner function ${\mathrm{Wig}}(v,\,u)$ of a pair $v,u$ of functions in ${\mathcal{S}}(\mathbb{R})$, and the map $\Psi$ establishes a one-to-one correspondence between tempered distributions in the plane and weakly continuous linear operators from ${\mathcal{S}}(\mathbb{R})$ to ${\mathcal{S}}^{\prime}(\mathbb{R})$. In particular, a collection of distributions $\mathfrak{S}$} is a weakly bounded subset of ${\mathcal{S}}^{\prime}(\mathbb{R}^{2})$ if and only if, for every pair $v,u$ of functions in ${\mathcal{S}}(\mathbb{R})$, the collection of scalar products $(v\,|\,\Psi({\mathfrak{S}})\,u)$ is bounded.
Setting $G=SL(2,\mathbb{R})$, the metaplectic group $\widetilde{G}$ is the (connected) twofold cover of $G$, and the metaplectic representation is a unitary representation $\mathrm{Met}$ of $\widetilde{G}$ in $L^{2}(\mathbb{R})$, the definition of which will be (almost) defined presently. If $g_{1}$ and $g_{2}$ are the two distinct points of $\widetilde{G}$ lying above the same point $g$ of $G$, one has ${\mathrm{Met}}(g_{1})=-{\mathrm{Met}}(g_{2})$, and we shall satisfy ourselves with defining ${\mathrm{Met}}(g_{1})$ up to the possible multiplication of $\pm 1$, so that it should depend only on $g$. The metaplectic representation is defined on generators of $G$ as follows. If $g=\left(\begin{smallmatrix}a&0\\ 0&a^{-1}\end{smallmatrix}\right)$ with $a>0$, ${\mathrm{Met}}(g)$ is (plus or minus) the operator $u\mapsto v$, with $v(x)=a^{-\frac{1}{2}}u(a^{-1}x)$; if $g=\left(\begin{smallmatrix}1&0\\ c&1\end{smallmatrix}\right)$, the same holds with $v(x)=u(x)\,e^{i\pi cx^{2}}$; if $g=\left(\begin{smallmatrix}0&1\\ -1&0\end{smallmatrix}\right)$, the operator ${\mathrm{Met}}(g)$ is $u\mapsto e^{-\frac{i\pi}{4}}\widetilde{u}$ with $\widetilde{u}(x)=2^{-\frac{1}{2}}\,\widehat{u}(\frac{x}{2})$. The rescaling by the factor $\frac{1}{2}$ is imposed by the covariance property to follow, and our choice of normalization of the Weyl calculus (justified in [21, section 6.1]).
The so-called covariance of the Weyl calculus under the metaplectic representation is expressed by any of the two equivalent identities, valid for every ${\mathfrak{S}}\in{\mathcal{S}}^{\prime}(\mathbb{R}^{2})$ and $v,u\in{\mathcal{S}}(\mathbb{R})$ :
${\mathrm{Met}}(g)\,\Psi({\mathfrak{S}})\,({\mathrm{Met}}(g))^{-1}$ $=\Psi\left({\mathfrak{S}}\,\circ\,g^{-1}\right),$
${\mathrm{Wig}}\,({\mathrm{Met}}(g)\,v,\,{\mathrm{Met}}(g)\,u)$ $={\mathrm{Wig}}(v,\,u)\,\circ\,g^{-1}.$ (7.3)
All this is very classical. The following fact seems to be less well-known. Borrowing from the Heisenberg representation the notation $P,Q$, where $P=\frac{1}{2i\pi}\,\frac{d}{dx}$ and $Q$ is the operator that multiplies a function of $x$ by $x$, one has the identity [20, p.55]
$P\,\Psi({\mathfrak{S}})\,Q-Q\,\Psi({\mathfrak{S}})\,P=\Psi({\mathcal{E}}\,{\mathfrak{S}}).$ (7.4)
This is a consequence of the simple formulas [19, p.27] which give, for any operator $A=\Psi({\mathfrak{S}})$, the symbols of the operators $PA,AP,QA$ and $AQ$.
The series ${\mathfrak{B}}=\frac{1}{2}\sum_{g\in\Gamma/\Gamma_{\infty}}{\mathfrak{s}}_{1}^{1}\,\circ\,g^{-1}$ does not quite define a distribution. It will become one after one has applied it the operator $(2i\pi{\mathcal{E}})^{2}$ (which may be applied instead to the Wigner function the object $\mathfrak{B}$} is tested on).
The general term of the series that defines $\mathfrak{B}$} is, with $g=\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right)$,
${\mathfrak{B}}_{a,c}(x,\xi)=\left({\mathfrak{s}}_{1}^{1}\,\circ\,g^{-1}\right)(x,\xi)=e^{2i\pi(dx-b\xi)}\delta(-cx+a\xi-1).$ (7.5)
That it depends only on the pair $a,c$ can also be checked from the fact that, given $\left(\begin{smallmatrix}a\\ c\end{smallmatrix}\right)$, the column $\left(\begin{smallmatrix}b\\ d\end{smallmatrix}\right)$ is unique up to the addition of a multiple of $\left(\begin{smallmatrix}a\\ c\end{smallmatrix}\right)$. When $a\neq 0$,
$\left({\mathfrak{s}}_{1}^{1}\,\circ\,g^{-1}\right)(x,\xi)=e^{-\frac{2i\pi b}{a}}\,e^{\frac{2i\pi x}{a}}\delta(-cx+a\xi-1).$ (7.6)
Dropping the constant factor of absolute value $1$ which does not change the estimates, one sets
$\overset{\circ}{{\mathfrak{B}}}_{a,c}(x,\xi)=e^{\frac{2i\pi x}{a}}\delta(-cx+a\xi-1).$ (7.7)
Theorem 7.1.
When $a\neq 0$, one has for $v,u\in{\mathcal{S}}(\mathbb{R})$ the identity
$\left(v\,\big|\,\Psi(\overset{\circ}{{\mathfrak{B}}}_{a,c})\,u\right)=|a|^{-1}\,A_{a,c}(v)\,B_{a,c}(u),$ (7.8)
with
$A_{a,c}(v)$ $=\frac{1}{2}\,\int_{-\infty}^{\infty}\overline{v}(x)\,e^{\frac{i\pi cx^{2}}{2a}}\,e^{\frac{2i\pi x}{a}}\,dx,$
$B_{a,c}(u)$ $=\frac{1}{2}\,\int_{-\infty}^{\infty}u(y)\,e^{-\frac{i\pi cy^{2}}{2a}}\,dy.$ (7.9)
Next,
$\left(v\,\big|\,\Psi({\mathcal{E}}^{2}\,\overset{\circ}{\mathfrak{B}}_{a,c})\,u\right)\\ =\frac{1}{4}\,|a|^{-3}e^{-\frac{2i\pi b}{a}}\,\int_{-\infty}^{\infty}\overline{v}(x)\,e^{\frac{i\pi\,cx^{2}}{2a}}\,e^{\frac{2i\pi x}{a}}\,dx\int_{-\infty}^{\infty}y^{2}\,u(y)\,e^{-\frac{i\pi\,cy^{2}}{2a}}\,dy.$ (7.10)
When $c\neq 0$, one has also (recall that $\widetilde{v}(x)=2^{-\frac{1}{2}}\widehat{v}(\frac{x}{2})$)
$\left(v\,\big|\,\Psi({\mathcal{E}}^{2}\,\overset{\circ}{\mathfrak{B}}_{a,c})\,u\right)\\ =\frac{1}{4}\,|c|^{-3}e^{\frac{2i\pi d}{c}}\,\int_{-\infty}^{\infty}\overline{\widetilde{v}(\xi)}\,e^{-\frac{i\pi\,a\xi^{2}}{2c}}\,e^{\frac{2i\pi\xi}{c}}\,d\xi\int_{-\infty}^{\infty}\eta^{2}\,\widetilde{u}(\eta)\,e^{\frac{i\pi\,a\eta^{2}}{2c}}\,d\eta.$ (7.11)
Proof.
The equation (7.2) provides the integral kernel $K_{a,c}$ of the operator $\Psi(\overset{\circ}{{\mathfrak{B}}}_{a,c})$. One has
$({\mathcal{F}}_{2}^{-1}\overset{\circ}{{\mathfrak{B}}}_{a,c})(x,\,z)=|a|^{-1}\,e^{\frac{2i\pi x}{a}}\,\exp\left(2i\pi z\,\frac{cx+1}{a}\right),$ (7.12)
and
$K_{a,c}(x,\,y)=({\mathcal{F}}_{2}^{-1}\overset{\circ}{{\mathfrak{B}}}_{a,c})\left(\frac{x+y}{2},\,\frac{x-y}{2}\right)=|a|^{-1}\,e^{\frac{2i\pi x}{a}}\,\exp\left(\frac{i\pi c(x^{2}-y^{2})}{2a}\right).$ (7.13)
The equation (7.8) follows.
The equation (7.4) gives
$\left(v\,\big|\,\Psi({\mathcal{E}}\,\overset{\circ}{{\mathfrak{B}}}_{a,c})\,u\right)=|a|^{-1}\,\left[A_{a,c}(Pv)\,B_{a,c}(Qu)-A_{a,c}(Q\overline{v})\,B_{a,c}(Pu)\right]$ (7.14)
After integrations by parts, one has
$A_{a,c}(Pv)$ $=-\frac{1}{4i\pi}\,\int_{-\infty}^{\infty}\overline{v}(x)\,e^{\frac{i\pi cx^{2}}{2a}}\,e^{\frac{2i\pi x}{a}}\left(\frac{i\pi cx}{a}+\frac{2i\pi}{a}\right)dx,$
$B_{a,c}(Pu)$ $=-\frac{1}{4i\pi}\,\int_{-\infty}^{\infty}u(y)\,e^{-\frac{i\pi cy^{2}}{a}}\,\frac{i\pi cy}{a}\,dy.$ (7.15)
In the difference (7.14), the terms involving the factor $xy$ under the integral sign cancel off. What remains gives
$\left(v\,\big|\,\Psi({\mathcal{E}}\,\overset{\circ}{{\mathfrak{B}}}_{a,c})\,u\right)=-\frac{|a|^{-1}}{2a}\,\left(v\,\big|\,\Psi(\overset{\circ}{{\mathfrak{B}}}_{a,c})\,(yu)\right).$ (7.16)
Making the same operation a second time, one obtains
$\left(v\,\big|\,\Psi({\mathcal{E}}^{2}\,\overset{\circ}{{\mathfrak{B}}}_{a,c})\,u\right)=\frac{1}{4\,|a|^{3}}\,\left(v\,\big|\,\Psi(\overset{\circ}{{\mathfrak{B}}}_{a,c})\,(y^{2}u)\right),$ (7.17)
which is just (7.6).
Next, if $g=\left(\begin{smallmatrix}a&\centerdot\\ c&\centerdot\end{smallmatrix}\right)$ and $g^{\prime}=\left(\begin{smallmatrix}0&1\\ -1&0\end{smallmatrix}\right)\,g=\left(\begin{smallmatrix}c&\centerdot\\ -a&\centerdot\end{smallmatrix}\right)$, one has according to the covariance rule (7)
$\Psi({\mathfrak{s}}_{1}^{1}\,\circ\,g^{-1})=\Psi\left({\mathfrak{s}}_{1}^{1}\,\circ\,g^{\prime-1}\,\circ\,\left(\begin{smallmatrix}0&1\\ -1&0\end{smallmatrix}\right)\right)={\mathcal{F}}_{\frac{1}{2}}\,\Psi({\mathfrak{s}}_{1}^{1}\,\circ\,g^{\prime-1})\,{\mathcal{F}}_{\frac{1}{2}}^{-1},$ (7.18)
with ${\mathcal{F}}_{\frac{1}{2}}\,u=\widetilde{u}$. The equation (7.11) is thus a consequence of (7.10).
Corollary 7.2.
The series (6.4) is convergent if $j\geq 1$, and defines ${\mathcal{E}}^{2}{\mathfrak{B}}$ as an element of ${\mathcal{S}}^{\prime}(\mathbb{R}^{2})$.
Proof.
Denoting as $\|\,\|_{\overset{}{1}}$ the norm in the space $L^{1}(\mathbb{R})$, we shall use if $ac\neq 0$, from (7.6) and (7.11), taking $u=u(y)$ and $\widetilde{u}=\widetilde{u}(\eta)$ the estimates
$|\left(v\,\big|\,\Psi((2i\pi{\mathcal{E}})^{2}{\mathfrak{B}}_{a,c})\,u\right)|\leq\begin{cases}|a|^{-3}|c|\,\times\,\pi^{2}\,(\|\,v\,\|_{\overset{}{1}}\,\|\,y^{2}u\,\|_{\overset{}{1}}\qquad&{\mathrm{when}}\,\,a\neq 0,\\ |c|^{-3}|a|\,\times\,\pi^{2}\,\|\,\widetilde{v}\,\|_{\overset{}{1}}\,\|\,\eta^{2}\widetilde{u}\,\|_{\overset{}{1}}\qquad&{\mathrm{when}}\,\,c\neq 0.\end{cases}$ (7.19)
This is the general term of a summable $(a,c)$-series. One concludes from the fact, recalled in the beginning of the present section, that weakly continuous linear operators form ${\mathcal{S}}(\mathbb{R})$ to ${\mathcal{S}}^{\prime}(\mathbb{R})$ are precisely the operators $\Psi({\mathfrak{S}})$ with ${\mathfrak{S}}\in{\mathcal{S}}^{\prime}(\mathbb{R}^{2})$.
The following two identities are facts of covariance (7).
Lemma 7.3.
Given a tempered distribution $\mathfrak{S}$} and $v,u\in{\mathcal{S}}(\mathbb{R})$, one has for every $q>0$ the identity
$\left(v\,\big|\,\Psi\left(q^{1-2i\pi{\mathcal{E}}^{\natural}}\,{\mathfrak{S}}\right)\,u\right)=\left(q\,v_{q}\,\big|\,\Psi\left({\mathfrak{S}}\right)\,(q\,u_{q})\right),$ (7.20)
with $v_{q}(x)=v(qx)$ and $u_{q}(y)=u(qy)$. Next, denoting as ${\mathcal{F}}_{\frac{1}{2}}$ the transformation $u\mapsto\widetilde{u}$, one has for every $\gamma\in\mathbb{R}$ the identity (cf. Lemma 5.2 for the definition of $\tau[\gamma]$)
$\tau[\gamma]\,{\mathrm{Wig}}(v,\,u)$ $={\mathrm{Wig}}\left({\mathrm{Met}}(\left(\begin{smallmatrix}1&-\gamma\\ 0&1\end{smallmatrix}\right))\,v,\,\,{\mathrm{Met}}(\left(\begin{smallmatrix}1&-\gamma\\ 0&1\end{smallmatrix}\right))\,v\right)$
[…]
Recall that ${\mathrm{Met}}(\left(\begin{smallmatrix}1&0\\ \gamma&1\end{smallmatrix}\right))$ is the operator that multiplies a function of $\xi$ by $e^{i\pi\gamma\xi^{2}}$.
Proof.
The first identity is a direct application of the definition of the metaplectic transformation: one has $\left({\mathrm{Met}}\left(\left(\begin{smallmatrix}q^{-1}&0\\ 0&q\end{smallmatrix}\right)\right)u\right)(x)=q^{\frac{1}{2}}u(qx)$. To obtain the second one, we must also use the observation that $\left(\begin{smallmatrix}1&-\gamma\\ 0&1\end{smallmatrix}\right)=\left(\begin{smallmatrix}0&-1\\ 1&0\end{smallmatrix}\right)\,\left(\begin{smallmatrix}1&0\\ \gamma&1\end{smallmatrix}\right)\,\left(\begin{smallmatrix}0&1\\ -1&0\end{smallmatrix}\right)$.
Piecing (7.20) and (7.3), one obtains, not forgetting (3.5),
$\left(v\,\big|\,\Psi\left(\tau_{-\gamma}\,\,q^{1-2i\pi{\mathcal{E}}^{\natural}}\,{\mathfrak{S}}\right)\,u\right)=<q^{1-2i\pi{\mathcal{E}}^{\natural}}\,{\mathfrak{S}},\,\,\tau[\gamma]\,{\mathrm{Wig}}(v,\,u)>\\ =<q^{1-2i\pi{\mathcal{E}}^{\natural}}\,{\mathfrak{S}}\,\,,\,\,{\mathrm{Wig}}\left({\mathrm{Met}}(\left(\begin{smallmatrix}1&-\gamma\\ 0&1\end{smallmatrix}\right))\,v,\,{\mathrm{Met}}(\left(\begin{smallmatrix}1&-\gamma\\ 0&1\end{smallmatrix}\right))\,u>\right)\\ =\left(\,q\,\left[{\mathrm{Met}}(\left(\begin{smallmatrix}1&-\gamma\\ 0&1\end{smallmatrix}\right))\,v\right]_{q}\,\,\big|\,\,\Psi({\mathfrak{S}})\,\left(\,q\,\left[{\mathrm{Met}}(\left(\begin{smallmatrix}1&-\gamma\\ 0&1\end{smallmatrix}\right))\,u\right]_{q}\right)\right)\\ \left(\,q\,\left[{\mathcal{F}}_{\frac{1}{2}}^{-1}({\mathrm{Met}}(\left(\begin{smallmatrix}1&0\\ \gamma&1\end{smallmatrix}\right))\,\widetilde{v})\right]_{q}\,\,\big|\,\,\Psi({\mathfrak{S}})\,\left(\,q\,\left[{\mathcal{F}}_{\frac{1}{2}}^{-1}({\mathrm{Met}}(\left(\begin{smallmatrix}1&0\\ \gamma&1\end{smallmatrix}\right))\,\widetilde{u})\right)\right]_{q}\right)$ (7.22)
The main role of pseudodifferential operators in the present proof is based on the fact that the Hecke operator $T_{p}^{\mathrm{dist}}$ and the operators $R^{k-2\ell}\,\sigma_{r}$ which make up the series (5) have a nice realization if applied to Wigner functions. The operator $2i\pi{\mathcal{E}}$ commutes with the action of $SL(2,\mathbb{R})$ on $\mathbb{R}^{2}$, hence with all operators concerned here ; the operator $(2i\pi{\mathcal{E}})^{2}$ can be moved from one side to the other of the scalar products to be considered.
Theorem 7.4.
Set ${\mathfrak{B}}^{[2]}={\mathcal{E}}^{2}\,{\mathfrak{B}}$ (a multiple of the distribution ${\mathfrak{B}}^{1}$ as so denoted in (6.4)). Given a prime $p$, and $q=p^{n}$ with $n\in\mathbb{Z}$, the distribution $q^{1-2i\pi{\mathcal{E}}^{\natural}}\,\sigma_{r}\,{\mathfrak{B}}^{[2]}$ remains in a weakly bounded subset of ${\mathcal{S}}^{\prime}(\mathbb{R}^{2})$ independent of $n$ and of  $r=0,1,\dots,n$.
Proof.
Recall that it suffices to bound the result of testing the distribution under consideration on ${\mathrm{Wig}}(v,\,u)$ with $v,u\in{\mathcal{S}}(\mathbb{R})$.
The distribution ${\mathfrak{B}}^{[2]}$ is invariant under the map $(x,\xi)\mapsto(x+\xi,\xi)$. Hence, it belongs to the space $\rm{Inv}$ as introduced immediately before Lemma 5.2. One can then apply (5.10) to move the factor $\sigma_{r}$ from the right to the left of $q^{1-2i\pi{\mathcal{E}}^{\natural}}$.
With the notation $R=p^{-\frac{1}{2}+i\pi{\mathcal{E}}^{\natural}}$, one has $q^{1-2i\pi{\mathcal{E}}^{\natural}}=R^{-2n}$ and we recall the identity
$R^{-2n}\sigma_{r}\sim\sigma_{-2n+r}R^{-2n}\quad{\mathrm{if}}\,\,n\leq 0,\qquad{\mathrm{and}}\qquad R^{-2n}\sigma_{r}\sim\sigma_{r}^{(2n)}\,R^{-2n}\quad{\mathrm{if}}\,\,n\geq 0.$ (7.23)
One obtains in the first case, i.e., when $q\leq 1$,
$q^{1-2i\pi{\mathcal{E}}^{\natural}}\,\sigma_{r}\,{\mathfrak{B}}^{[2]},\,{\mathrm{Wig}}(v,u)>=<q^{1-2i\pi{\mathcal{E}}^{\natural}}\,{\mathfrak{B}}^{[2]},\,\,\sigma_{-2n+r}\,{\mathrm{Wig}}(v,u)>\\ =<q^{1-2i\pi{\mathcal{E}}^{\natural}}\,{\mathfrak{B}}^{[2]},\,\,\,p^{2n-r}\sum_{0\leq b<p^{r-2n}}\tau\left(\frac{b}{p^{2n-r}}\right)\,{\mathrm{Wig}}(v,u)>,$ (7.24)
an arithmetic average of expressions $q^{1-2i\pi{\mathcal{E}}^{\natural}}\,{\mathfrak{B}}^{[2]},\,\,\,\tau[\gamma]\,{\mathrm{Wig}}(v,u)>$ with $0\leq\gamma<1$.
Using Lemma 7.3, such an expression is the same as
[…]
Now, one has
$\left({\mathcal{F}}_{\frac{1}{2}}^{-1}\,(e^{i\pi\gamma\xi^{2}}\,\widetilde{u})\right)(x)=2^{\frac{1}{2}}\int_{-\infty}^{\infty}e^{4i\pi x\xi}e^{i\pi\gamma\xi^{2}}\,\widetilde{u}(\xi)\,d\xi.$ (7.26)
This function of $x$ is bounded in a uniform way relative to $\gamma$ and, as seen with the help of an integration by parts, so is its product by $x^{2}$ since $\gamma$ is bounded. The $L^{1}$-norm of this function is thus bounded and, applying Corollary 7.2, we are done for this case.
This proof does not extend to the case when $n\geq 0$, i.e., $q>1$, since the transform
$\sigma_{r}^{(2n)}=p^{-r}\sum_{0\leq b<p^{r}}\tau[bp^{2n-r}]$ (7.27)
involves then translations $\tau$ with unbounded amplitudes: observe however that $\gamma=bp^{2n-r}\leq p^{2n}=q^{2}$.
The effect of the transformation ${\mathrm{Met}}\left(\left(\begin{smallmatrix}1&-\gamma\\ 0&1\end{smallmatrix}\right)\right)$ is to multiply the ${\mathcal{F}}_{\frac{1}{2}}$-transform of the function it is applied to by $e^{i\pi\gamma\eta^{2}}$. We use the equation (7.11) when $|c|\geq|a|$: it is then immediate that the introduction of such a factor does not destroy the second case of the estimate (7.19).
The last case is that for which $|c|\leq|a|$ : dispensing with the trivial case for which $c=0,a=\pm 1$, we use then (7.10) and (7.22) with ${\mathfrak{S}}=\overset{\circ}{I}_{a,c}$. Specializing in the second factor (the first is totally similar), we write
$B_{a,c}\left(q\,\left[{\mathcal{F}}_{\frac{1}{2}}^{-1}\left({\mathrm{Met}}(\left(\begin{smallmatrix}1&0\\ \gamma&1\end{smallmatrix}\right))\,\widetilde{u})\right)\right]_{q}\right)\\ =2^{-\frac{1}{2}}q\,\int_{-\infty}^{\infty}e^{-\frac{i\pi cy^{2}}{2a}}\,dy\int_{-\infty}^{\infty}\widetilde{u}(\eta)\,e^{i\pi\gamma\eta^{2}}\,e^{4i\pi qy\eta}\,d\eta.$ (7.28)
Now,
$\int_{-\infty}^{\infty}e^{-\frac{i\pi cy^{2}}{2a}}\,e^{4i\pi qy\eta}\,dy=\left(\frac{ic}{2a}\right)^{-\frac{1}{2}}\,\exp\left(\frac{8i\pi a\,q^{2}\eta^{2}}{c}\right),$ (7.29)
and
$B_{a,c}\left(q\,\left[{\mathcal{F}}_{\frac{1}{2}}^{-1}\left({\mathrm{Met}}(\left(\begin{smallmatrix}1&0\\ \gamma&1\end{smallmatrix}\right))\,\widetilde{u})\right)\right]_{q}\right)=\left(\frac{ic}{a}\right)^{-\frac{1}{2}}\,\int_{-\infty}^{\infty}\widetilde{u}(\eta)\,\exp\left[i\pi(\gamma+\frac{8a\,q^{2}}{c})\,\eta^{2}\right]\,d\eta.$ (7.30)
As seen in (7.10), using ${\mathcal{E}}^{2}\,\overset{\circ}{\mathfrak{B}}$ in place of $\overset{\circ}{\mathfrak{B}}$, we benefit from an extra factor $\eta^{2}$ in the last integral. Since, for $|c|\leq|a|$, $|(\gamma+\frac{8a\,q^{2}}{c})\,\eta^{2}|\geq|\frac{4aq^{2}\eta^{2}}{c}|$, an integration by parts using the identity
$\eta\,\exp\left[i\pi(\gamma+\frac{8a\,q^{2}}{c})\,\eta^{2}\right]=\left(2i\pi(\gamma+\frac{8a\,q^{2}}{c})\right)^{-1}\,\frac{d}{d\eta}\,\exp\left[i\pi(\gamma+\frac{8a\,q^{2}}{c})\,\eta^{2}\right]$ (7.31)
makes it possible to obtain for $B_{a,c}\left(q\,\left[{\mathcal{F}}_{\frac{1}{2}}^{-1}\left({\mathrm{Met}}(\left(\begin{smallmatrix}1&0\\ \gamma&1\end{smallmatrix}\right))\,\widetilde{u})\right)\right]_{q}\right)$ bounds by arbitrary powers of $q^{-1}$.
The summability with respect to the pair $a,c$ is ensured by the factor $|a|^{-3}$ in front of (7.10).
8. The Ramanujan-Petersson estimate for Maass forms
Lemma 8.1.
Given a finite collection $\left({\mathfrak{N}}_{j}\right)_{j\geq 1}$ of distinct Hecke distributions, one can find $h$ in ${\mathcal{S}}(\mathbb{R}^{2})$ such that $\big\langle{\mathfrak{N}}_{1},h\big\rangle\,=1$ but $\big\langle{\mathfrak{N}}_{j},h\big\rangle\,=0$ for $j>1$.
Proof.
Let $\mathfrak{T}$} be the linear form on the linear space generated by the ${\mathfrak{N}}_{j}$ defined by the conditions $\big\langle T,{\mathfrak{N}}_{1}\big\rangle\,=1$ but $\big\langle T,{\mathfrak{N}}_{j}\big\rangle\,=0$ for $j>1$. The Hahn-Banach theorem makes it possible to extend $T$ as a continuous linear form on ${\mathcal{S}}^{\prime}(\mathbb{R}^{2})$. But the space ${\mathcal{S}}(\mathbb{R}^{2})$ is reflexive, which means that this extension is associated in the natural way with a function $h\in{\mathcal{S}}(\mathbb{R}^{2})$.
Theorem 8.2.
Let a prime $p$ and $\varepsilon>0$ be given. As $N\to\infty$, the (even) distribution $2^{-2N}\left(T_{p}^{\mathrm{dist}}\right)^{2N}\,P_{2}(-2i\pi{\mathcal{E}})\,{\mathfrak{B}}$ remains in a bounded subset of ${\mathcal{S}}^{\prime}(\mathbb{R}^{2})$.
Proof.
Recall the decomposition (5.20) of the distribution $\left(T_{p}^{\mathrm{dist}}\right)^{2N}{\mathfrak{B}}$, and let $\varepsilon>0$ be given. From Theorem 7.4, one has for some continuous norm $|\big\bracevert\,\,\big\bracevert$ on ${\mathcal{S}}(\mathbb{R}^{2})$ the estimates
$|,\left|\,\big\langle p^{(N-\ell)(-1+2i\pi{\mathcal{E}}^{\natural})}\,\sigma_{r}\,P_{2}(-2i\pi{\mathcal{E}})\,{\mathfrak{B}},\,h\big\rangle\,\right|\leq\big\bracevert\,h\,\big\bracevert$, (8.1)
valid whenever $h\in{\mathcal{S}}(\mathbb{R}^{2})$. To avoid having to carry the operator $P_{2}(-2i\pi{\mathcal{E}})$, we set $g=P_{2}(2i\pi{\mathcal{E}})\,h$. We shall move $P_{2}(2i\pi{\mathcal{E}})$ to the other side later, not forgetting that $2i\pi{\mathcal{E}}$ commutes with all operators involved. The effect of applying $P_{2}(-2i\pi{\mathcal{E}})$ to a Hecke distribution ${\mathfrak{N}}_{\chi,i\lambda}$ is to multiply it by $P_{2}(i\lambda)$
Adding these inequalities and recalling from Proposition 5.3 that, for any given $\ell$, one has $\alpha_{2N,\ell}^{(0)}+\alpha_{2N,\ell}^{(1)}+\dots+\alpha_{2N,\ell}^{(\ell)}=\left(\begin{smallmatrix}2N\\ \ell\end{smallmatrix}\right)$, we obtain with $g=P_{2}(2i\pi{\mathcal{E}})\,h$
$\sum_{\ell=0}^{2N}\sum_{0\leq r\leq\ell}\alpha_{2N,\ell}^{(r)}\,$ $|\displaystyle\left|\,\big\langle p^{(N-\ell)(-1+2i\pi{\mathcal{E}}^{\natural})}\,\sigma_{r}{\mathfrak{B}},\,g\big\rangle\,\right|\leq\sum_{\ell=0}^{2N}\sum_{0\leq r\leq\ell}\alpha_{2N,\ell}^{(r)}\,\big\bracevert h\big\bracevert$
[…]
This is the claimed result.
We remark now that Theorem 8.2 remains valid if we replace $\mathfrak{B}$} by either term of the decomposition ${\mathfrak{B}}={\mathfrak{B}}^{\mathrm{cont}}+{\mathfrak{B}}^{\mathrm{disc}}$, as defined in reference to the continuous and discrete parts of the spectral decomposition (6.5).
Proposition 8.3.
The statement of Theorem 8.2 remains valid if one substitutes ${\mathfrak{B}}^{\mathrm{cont}}$ for $\mathfrak{B}$}.
Proof.
The equation (5.7) reduces in the case of the Eisenstein distribution ${\mathfrak{E}}_{i\lambda}$ to $T_{p}^{\mathrm{dist}}{\mathfrak{E}}_{i\lambda}=\left(p^{\frac{i\lambda}{2}}+p^{-\frac{i\lambda}{2}}\right){\mathfrak{E}}_{i\lambda}$. The insertion of the bounded factor $p^{\frac{i\lambda}{2}}+p^{-\frac{i\lambda}{2}}$ in the integrand of the integral part of (6.5) does not destroy its summability.
It is, for our present purpose, the difference ${\mathfrak{B}}^{\mathrm{disc}}={\mathfrak{B}}-{\mathfrak{B}}^{\mathrm{cont}}$ we are interested in. It follows from Theorem 8.2 and Proposition 8.3 that there exists a continuous semi-norm $|\big\bracevert\,\,\big\bracevert$ on ${\mathcal{S}}(\mathbb{R}^{2})$, independent of $N$, such that, for $h\in{\mathcal{S}}(\mathbb{R}^{2})$ and $g=P_{2}(2i\pi{\mathcal{E}})\,h$,
$|.\big|\,\,\big\langle\left(T_{p}^{\mathrm{dist}}\right)^{2N}{\mathfrak{B}}^{\mathrm{disc}},\,g\big\rangle\,\,\big|\leq 2^{2N}\,\big\bracevert h\big\bracevert$. (8.3)
The last point of this paper will be a localization (in the spectral sense, with respect to the discrete part of the spectrum of the automorphic Euler operator) of this estimate. To this effect, with a function $F$ such that, for every $B$, $|F(t)|\leq C\,e^{-B\,|t|}$ for some $C>0$, introduce the function
$\Phi(i\lambda)=\int_{-\infty}^{\infty}F(t)\,e^{-2i\pi t\lambda}dt=\frac{1}{2\pi}\int_{0}^{\infty}F\left(\frac{\log\theta}{2\pi}\right)\theta^{-i\lambda-1}d\theta.$ (8.4)
One has
$\frac{1}{2\pi}\int_{0}^{\infty}\left|F\left(\frac{\log\theta}{2\pi}\right)\right|\,\frac{d\theta}{\theta}=\int_{-\infty}^{\infty}|F(t)|\,dt<\infty.$ (8.5)
Since the operators $\theta^{2i\pi{\mathcal{E}}}$ make up a collection of unitary operators, one can substitute $2i\pi{\mathcal{E}}$ for $i\lambda$ in (8.4), getting the definition of a bounded operator in $L^{2}(\mathbb{R}^{2})$, to wit the operator $\Phi(-2i\pi{\mathcal{E}})$ such that, for $g\in L^{2}(\mathbb{R}^{2})$,
$(\Phi(-2i\pi{\mathcal{E}})\,g)(x,\xi)=\int_{-\infty}^{\infty}F(t)\,g\left(e^{2\pi t}x,e^{2\pi t}\xi\right)e^{2\pi t}dt.$ (8.6)
One observes that if $g$ lies in ${\mathcal{S}}(\mathbb{R}^{2})$, so does $\Phi(-2i\pi{\mathcal{E}})\,g$, and one can define $\Phi(2i\pi{\mathcal{E}})\,{\mathfrak{B}}^{\mathrm{disc}}$ by duality and, making use of the second term on the right-hand side of (6.5), one obtains
$\,\big\langle\Phi(2i\pi{\mathcal{E}})\,{\mathfrak{B}}^{\mathrm{disc}},\,g\big\rangle\,$ $=\,\big\langle{\mathfrak{B}}^{\mathrm{disc}},\,\Phi(-2i\pi{\mathcal{E}})\,g\big\rangle\,$
$=\frac{1}{2}\sum_{\begin{array}[]{c}r,\iota\\ r\in\mathbb{Z}^{\times}\end{array}}\Phi(-i\lambda_{r})\,\frac{\Gamma(1-\frac{i\lambda_{r}}{2})\Gamma(1+\frac{i\lambda_{r}}{2})}{\|{\mathcal{N}}^{|r|,\iota}\|^{2}}\,\big\langle{\mathfrak{N}}^{r,\iota},\,g\big\rangle\,$, (8.9)
which leads in view of (4.28) to the unsurprising formula
$\big\langle\Phi(2i\pi{\mathcal{E}})\left(T_{p}^{\mathrm{dist}}\right)^{2N}{\mathfrak{B}}^{\mathrm{disc}},\,g\big\rangle\,\\ =\frac{1}{2}\sum_{\begin{array}[]{c}r,\iota\\ r\in\mathbb{Z}^{\times}\end{array}}\Phi(-i\lambda_{r})\,(b_{p}(r,\iota))^{2N}\frac{\Gamma(1-\frac{i\lambda_{r}}{2})\Gamma(1+\frac{i\lambda_{r}}{2})}{\|{\mathcal{N}}^{|r|,\iota}\|^{2}}\,\big\langle{\mathfrak{N}}^{r,\iota},\,g\big\rangle\,$, (8.10)
where $b_{p}(r,\iota)$ denotes the $p$th Fourier coefficient of the series (4.18) for ${\mathcal{N}}^{r,\iota}$. The series (8.10) is absolutely convergent as a consequence of Theorem 6.1, since the insertion of the bounded factor $\Phi(-i\lambda_{r})\,(b_{p}(r,\iota))^{2N}$ does not change this state of affairs.
To isolate an eigenvalue $\frac{1+\lambda_{r}^{2}}{4}$ of the automorphic Laplacian or at least, taking advantage of Lemma 8.1, the set of eigenvalues $\frac{1+\lambda_{s}^{2}}{4}$ in some finite interval containing the point $\frac{1+\lambda_{r}^{2}}{4}$, we take
$\Phi_{N}(-i\lambda)=\exp\left(-\pi N\beta(\lambda-\lambda_{r})^{2}\right),$ (8.11)
with a small positive constant $\beta$ to be chosen later, $N$ being the large integer already so denoted.
The topology of ${\mathcal{S}}(\mathbb{R}^{2})$ can be defined by a “directed” family of semi-norms of the kind $\sum_{j}\,{\mathrm{sup}}\left|P_{j}\left(x,\frac{\partial}{\partial x},\xi,\frac{\partial}{\partial\xi}\right)h\right|$, where the $P_{j}$’s make up a finite collection of ordered monomials in the non-commuting operators indicated, of degrees $\leq A$: such a semi-norm will be said to be of degree $\leq A$.
Proposition 8.4.
Assume that the semi-norm $|\big\bracevert h\big\bracevert$ present in (8.3) is of degree $\leq A$. Given $\beta$, one has for $h\in{\mathcal{S}}(\mathbb{R}^{2})$ and $g=P_{2}(2i\pi{\mathcal{E}})\,h$ the estimate
$||\big|\,\big\langle\Phi_{N}(2i\pi{\mathcal{E}})\left(T_{p}^{\mathrm{dist}}\right)^{2N}\,{\mathfrak{B}}^{\mathrm{disc}},\,g\big\rangle\,\big|\leq\,2^{2N}\,\exp(\pi(A+1)^{2}N\beta)\,\big\bracevert\!\!\!\big\bracevert h\big\bracevert\!\!\!\big\bracevert$ (8.12)
for some new continuous norm $||\big\bracevert\!\!\!\big\bracevert\,\,\big\bracevert\!\!\!$$\big\bracevert$ on ${\mathcal{S}}(\mathbb{R}^{2})$.
Proof.
Using (8.6), we write
$\big\langle\Phi_{N}(2i\pi{\mathcal{E}})\left(T_{p}^{\mathrm{dist}}\right)^{2N}{\mathfrak{B}}^{\mathrm{disc}},\,g\big\rangle\,=\,\big\langle\left(T_{p}^{\mathrm{dist}}\right)^{2N}{\mathfrak{B}}^{\mathrm{disc}},\,\Phi_{N}(-2i\pi{\mathcal{E}})\,g\big\rangle\,\,$, (8.13)
and $g_{1}=\Phi_{N}(-2i\pi{\mathcal{E}})\,g$ is given as
$g_{1}(x,\xi)=\int_{-\infty}^{\infty}F_{N}(t)\,g\left(e^{2\pi t}x,e^{2\pi t}\xi\right)e^{2\pi t}dt,$ (8.14)
with
$F_{N}(t)$ $=\int_{-\infty}^{\infty}e^{2i\pi\lambda t}\Phi_{N}(i\lambda)\,d\lambda=(N\beta)^{-\frac{1}{2}}\exp\left(-\frac{\pi t^{2}}{N\beta}\right)e^{2i\pi\lambda_{r}t}$
$=e^{2i\pi\lambda_{r}t}(N\beta)^{-\frac{1}{2}}\exp\left(-\frac{\pi t^{2}}{N\beta}\right).$ (8.15)
Since the operator $2i\pi{\mathcal{E}}$ commutes with $T_{p}^{\mathrm{dist}}$, the link between $g$ and $g_{1}$ provided by (8.14) connects also $h$ and $h_{1}=\Phi_{N}(-2i\pi{\mathcal{E}})\,h$. An application of any of the $4$ operators $x,\frac{\partial}{\partial x},\xi,\frac{\partial}{\partial\xi}$ to a function evaluated at $(e^{2\pi t}x,e^{2\pi t}\xi)$ may lead to a loss by a factor $e^{2\pi\,|t|}$ at the most, so that
[…]
for some new norm $||\big\bracevert\!\!\!\big\bracevert\,\,\big\bracevert\!\!\!$$\big\bracevert$, still of the kind recalled immediately after (8.11). The estimate (8.12) follows from this inequality, (8.13) and (8.3).
Theorem 8.5.
(the Ramanujan-Petersson conjecture) Given a pair $\chi,\lambda$, the character $\chi$ has to be unitary if the distribution ${\mathfrak{N}}_{\chi,i\lambda}$ is modular. In other words, given a Hecke eigenform ${\mathcal{N}}^{r,\iota}$ and a prime $p$, and setting $T_{p}\,{\mathcal{N}}^{r,\iota}=b_{p}(r,\iota)\,{\mathcal{N}}^{r,\iota}$, one has the inequality $|b_{p}(r,\iota)|\leq 2$.
Proof.
In what follows, the fixed number $A$ is the one (originating from (8.12)) present in Proposition 8.4. Our aim is to prove that, given $r,\iota$ and $\delta>1$, one must have $|b_{p}(r,\iota)|\leq 2\delta$. Assuming $|b_{p}(r,\iota)|\geq 2$, let $\alpha$ be such that
$|b_{p}(s,\kappa)|\leq\alpha\,|b_{p}(r,\iota)|$ (8.17)
for every $s,\kappa$. Next, we choose positive numbers $\beta,\,\eta$ in this order such that
$\exp\left(\frac{\pi(A+1)^{2}\beta}{2}\right)\leq\delta,\quad\alpha\,\exp\left(-\frac{\pi\beta\,\eta^{2}}{2}\right)\leq\frac{1}{4}.$ (8.18)
We note that there is only a finite number of eigenvalues $i\lambda_{s}$ of the automorphic Euler operator (including $i\lambda_{r}$) such that $|\lambda_{s}-\lambda_{r}|\leq\eta$. To single out ${\mathfrak{N}}^{r,\iota}$ within the finite collection of Hecke distributions ${\mathfrak{N}}^{s,\kappa}$ such that $|\lambda_{s}-\lambda_{r}|\leq\eta$, we use Lemma 8.1, obtaining a function $h\in{\mathcal{S}}(\mathbb{R}^{2})$ such that $\big\langle{\mathfrak{N}}^{r,\iota},h\big\rangle\,=1$ and $\big\langle{\mathfrak{N}}^{s,\kappa},h\big\rangle\,=0$ for $(s,\kappa)\neq(r,\iota)$ and $|\lambda_{s}-\lambda_{r}|\leq\eta$. Then, with $g=P_{2}(2i\pi{\mathcal{E}})\,h$, one has $\big\langle{\mathfrak{N}}^{r,\iota},g\big\rangle\,=\lambda_{r}^{2}(1+\lambda_{r}^{2})$.
We use the series (8.10), in which we substitute a general $(s,\kappa)$ for $(r,\iota)$, obtaining
$\big\langle\Phi(2i\pi{\mathcal{E}})\left(T_{p}^{\mathrm{dist}}\right)^{2N}{\mathfrak{B}}^{\mathrm{disc}},\,g\big\rangle\,=\sum_{\begin{array}[]{c}s,\kappa\\ s\in\mathbb{Z}^{\times}\end{array}}\Phi(-i\lambda_{s})\,(b_{p}(s,\kappa))^{2N}\,R_{N}(p,s,\kappa;\,g),$ (8.19)
with
$R_{N}(p,s,\kappa;\,g)=\frac{1}{2}\,\frac{\Gamma(1-\frac{i\lambda_{s}}{2})\Gamma(1+\frac{i\lambda_{s}}{2})}{\|{\mathcal{N}}^{s,\kappa}\|^{2}}\,\big\langle{\mathfrak{N}}^{s,\kappa},\,g\big\rangle\,.$ (8.20)
The term $R_{N}(p,s,\kappa;\,g)$ is the general term of the series on the right-hand side of (6.5), the absolute convergence of which was established in Proposition 6.1: the fact that $\Phi_{N}(-i\lambda_{s})\leq 1$ then proves the convergence of the series (8.19). We now analyze separately the sum of terms of this series for which $|\lambda_{s}-\lambda_{r}|>\eta$ or $|\lambda_{s}-\lambda_{r}|\leq\eta$. The point is to show that the second sum is dominant.
Using the facts that $|b_{p}(s,\kappa)|\leq\alpha\,|b_{p}(r,\iota)|$ and that, in the first case, $\Phi_{N}(-i\lambda_{s})\leq\exp\left(-\pi N\beta\,\eta^{2}\right)$, we obtain if using the last constraint (8.18)
$\Phi_{N}(-i\lambda_{s})\,|b_{p}(s,\kappa)|^{2N}$ $\leq\,e^{-\pi N\beta\eta^{2}}\alpha^{2N}|b_{p}(r,\iota)|^{2N}$
$\leq\left(\frac{1}{4\alpha}\right)^{2N}\alpha^{2N}|b_{p}(r,\iota)|^{2N}=4^{-2N}|b_{p}(r,\iota)|^{2N}.$ (8.21)
It thus follows from Proposition 6.1 that
[…]
For the pairs $(s,\kappa)$ with $|\lambda_{s}-\lambda_{r}|\leq\eta$, we recall the choice of $h$ made immediately after (8.18), from which it follows that the only contributing Hecke eigenform is ${\mathfrak{N}}^{r,\iota}$ and the corresponding term of the series (8.19) is (not forgetting that $P_{2}(i\lambda_{r})=\lambda_{r}^{2}(1+\lambda_{r}^{2})$)
$\Phi(-i\lambda_{r})\,(b_{p}(r,\iota))^{2N}\,R_{N}(p,r,\iota;\,g)=\frac{1}{2}\,(b_{p}(r,\iota))^{2N}\frac{\Gamma(1-\frac{i\lambda_{r}}{2})\Gamma(1+\frac{i\lambda_{r}}{2})}{\|{\mathcal{N}}^{r,\iota}\|^{2}}\,\,\big\langle{\mathfrak{N}}^{r,\iota},\,g\big\rangle\,\\ =\frac{1}{2}\,(b_{p}(r,\iota))^{2N}\frac{\Gamma(1-\frac{i\lambda_{r}}{2})\Gamma(1+\frac{i\lambda_{r}}{2})}{\|{\mathcal{N}}^{r,\iota}\|^{2}}\,\lambda_{r}^{2}(1+\lambda_{r}^{2}).$ (8.23)
Comparing the results of the last two equations, we see that, for $N$ large enough, the contribution of the series (8.22) is less than half that of the term (8.23) we are interested in. It thus follows from (8.12) that one half of the last expression (8.23) is at most
[…]
Letting $N\to\infty$, we are done.
9. The Selberg conjecture
We consider the Hecke subgroup $\Gamma_{0}(M)$ of $SL(2,\mathbb{Z})$ consisting of matrices $g=\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right)$ such that $c\equiv 0\,{\mathrm{mod}}\,M$, in order to benefit from the detailed structure of the space $\)L^{2}(\Gamma_{0}(M)\backslash\mathbb{H}$) as given in [3]. The Selberg conjecture is the fact that the spectrum of $\Delta$ in this Hilbert space coincides with $\frac{1}{4},\infty[$. It amounts to the fact that, given any Hecke eigenform $\mathcal{N}$}, an eigenfunction of $\Delta$ for the eigenvalue $\frac{1-\nu^{2}}{4}$, $\nu$ is pure imaginary: since this is a statement concerning individual Hecke distributions, one may assume, changing $M$ to a factor of $M$ if necessary, that $\mathcal{N}$} is a so-called new Hecke eigenform, so that it is still possible to normalize it by the fact that its first Fourier coefficient $b_{1}$, as characterized by the expansion [3, p.226]
${\mathcal{N}}(x+iy)=y^{\frac{1}{2}}\sum_{k\neq 0}b_{k}\,K_{\frac{\nu}{2}}(2\pi\,|k|\,y)\,e^{2i\pi kx},$ (9.1)
is $1$.
Such a modular form can be lifted to the distribution (4.19)
${\mathfrak{N}}(x,\xi)=\frac{1}{2}\sum_{k\neq 0}b_{k}\,\,|k|^{\frac{\nu}{2}}\,\,|\xi|^{-1-\nu}\exp\left(\frac{2i\pi kx}{\xi}\right),$ (9.2)
where $\nu$ is any of the two square roots of $\nu^{2}$ if ${\mathrm{Re\,}}(\nu^{2})<0$: we may assume that ${\mathrm{Re\,}}\nu>0$ if ${\mathrm{Re\,}}(\nu^{2})>0$. One has then $\Theta\,{\mathfrak{N}}={\mathcal{N}}$.
Fixing $M$ and $\Gamma=\Gamma_{0}(M)$, we find in [3, p.227] the Poincaré series, in which $j=1,2,\dots$,
$U_{1}\left(z,j+\frac{1}{2}\right)$ $=\frac{1}{2}\sum_{\tau\in\Gamma_{\infty}\backslash\Gamma}\left({\mathrm{Im\,}}(\tau\,.\,z)\right]^{j+\frac{1}{2}}\,\exp(2i\pi(\tau\,.\,z))$
$=\frac{1}{2}\sum_{\left(\begin{smallmatrix}n&n_{1}\\ m&m_{1}\end{smallmatrix}\right)\in\Gamma/\Gamma_{\infty}}\left(\frac{{\mathrm{Im\,}}z}{|-mz+n|^{2}}\right)^{j+\frac{1}{2}}\exp\left(2i\pi\frac{m_{1}z-n_{1}}{-mz+n}\right).$ (9.3)
We have specialized in the cusp $\infty$ of $\Gamma_{0}(M)\backslash\mathbb{H}$ and specialized to the value $1$ the coefficient in the exponent.
The function
$(4\pi)^{j}\frac{\Gamma(\frac{1}{2}+j)}{\Gamma(\frac{1}{2})}\,\times\,U_{1}\left(z,j+\frac{1}{2}\right)$ is the image under $\Theta$ of the distribution (of course with $\Gamma=\Gamma_{0}(M)$)
${\mathfrak{B}}^{j}=\frac{1}{2}\sum_{g\in\Gamma/\Gamma_{\infty}}\left({\mathfrak{s}}_{1}^{1}\right)^{j}\,\circ\,g^{-1}.$ (9.4)
Note that it was essential to start with ${\mathfrak{s}}_{1}^{1}$, not with ${\mathfrak{s}}_{1}^{a}$ (cf. (6.1)) for another value of $a\in\mathbb{Z}$, so that ${\mathfrak{B}}^{j}$ should be ${\mathcal{F}}^{\mathrm{symp}}$-invariant: this explains the necessity of taking for $\mathcal{N}$} a new Hecke eigenform, more precisely of choosing $M$ so that it should become one.
Now, the spectral decomposition of the Poincaré series $U_{1}\left(z,j+\frac{1}{2}\right)$ is given in [3, p.248] as the equation
$U_{1}\left(z,j+\frac{1}{2}\right)=\sum_{k}\frac{\left({\mathcal{N}}_{k}\,\big|\,U_{1}\left(\,\centerdot\,,j+\frac{1}{2}\right)\right)\,{\mathcal{N}}_{k}(z)}{\|\,{\mathcal{N}}_{k}\|^{2}}\\ +\frac{1}{4i\pi}\sum_{\mathfrak{c}}\int_{{\mathrm{Re\,}}\nu=0}\left(E_{\mathfrak{c}}(\,\centerdot\,,\,\frac{1-\nu}{2})\,\big|\,U_{1}\left(\,\centerdot\,,\,j+\frac{1}{2}\right)\right)\,E_{\mathfrak{c}}(\,\centerdot\,,\,\frac{1-\nu}{2})(z),$ (9.5)
the ingredients of which are as follows. The functions ${\mathcal{N}}_{k}$ are a complete set of new Hecke eigenforms, normalized in Hecke’s way, the letter $\mathfrak{c}$} runs through a complete set of inequivalent cusps, and $E_{\mathfrak{c}}(\,\centerdot\,,\,\frac{1-\nu}{2})$ is a non-holomorphic modular form (not a cusp-form) of Eisenstein type. There is no need to be more explicit regarding the Eisenstein terms since the integral term on the right-hand side of (9.5) is the spectral projection of $U_{1}\left(\,\centerdot\,,\,j+\frac{1}{2}\right)$ corresponding to the continuous part of the spectrum of $\Delta$.
Proposition 9.1.
Given $j=1,2,\dots$, the $\Gamma$-automorphic distribution $q^{2i\pi{\mathcal{E}}}{\mathfrak{B}}^{j}$ remains for $q\in]0,\infty[$ in a bounded subset of ${\mathcal{S}}^{\prime}(\mathbb{R}^{2})$.
Proof.
This proof is much simpler than that of Theorem 7.4, and one may dispense with pseudodifferential analysis here. As $SL(2,\mathbb{Z})$ has been replaced by $\Gamma=\Gamma_{0}(M)$, the class in $\Gamma/\Gamma_{\infty}$ of the matrix $\left(\begin{smallmatrix}n&n_{1}\\ m&m_{1}\end{smallmatrix}\right)$ is characterized by the pair $n,m$ with $m\equiv 0\,{\mathrm{mod}}\,M$. One has
$\big\langle{\mathfrak{B}},\,h\big\rangle\,=\frac{1}{2}\sum_{\begin{array}[]{c}m,n\in\mathbb{Z}\\ m\equiv 0\,{\mathrm{mod}}\,M\end{array}}I_{n,m}(h),$ (9.6)
with
$I_{n,m}(h)=\int_{-\infty}^{\infty}h(nx+n_{1},\,mx+m_{1})\,e^{2i\pi x}\,dx.$ (9.7)
Dropping, as may be done in view of the desired estimates, a constant factor of absolute value $1$, we write
$\overset{\circ}{I}_{n,m}\left(q^{2i\pi{\mathcal{E}}}h\right)$ $=q\int_{-\infty}^{\infty}h\left(q(nx-\frac{1}{m}\right),\,qmx)\,e^{2i\pi x}\,dx$
$=\int_{-\infty}^{\infty}h(nx-\frac{q}{m},\,mx)\,\exp\left(\frac{2i\pi x}{q}\right)dx.$ (9.8)
We can also write instead, using (3.4),
$\overset{\circ}{I}_{n,m}\left(q^{2i\pi{\mathcal{E}}}h\right)$ $=\int_{\mathbb{R}^{3}}\left({\mathcal{F}}^{\mathrm{symp}}h\right)(y,\eta)\,\exp\left(\frac{2i\pi x}{q}\right)$
$\exp\left[2i\pi\left((nx-\frac{q}{m})\eta-ymx\right)\right]dx\,dy\,d\eta.$ (9.9)
Since, for $m\neq 0$,
$\int_{-\infty}^{\infty}\exp\left(2i\pi(n-my+\frac{1}{q})\,x\right)dx=\delta(n\eta-my+\frac{1}{q})=\frac{1}{|m|}\delta\left(y-\frac{n\eta}{m}-\frac{1}{qm}\right),$ (9.10)
one obtains if $m\neq 0$
$\overset{\circ}{I}_{n,m}\left(q^{2i\pi{\mathcal{E}}}h\right)=\frac{1}{|m|}\int_{-\infty}^{\infty}\left({\mathcal{F}}^{\mathrm{symp}}h\right)\left(\frac{n\eta}{m}+\frac{1}{qm},\,\eta\right)\,d\eta.$ (9.11)
We write
$\left(\frac{n\eta}{m}+\frac{1}{qm}\right)^{2}+\eta^{2}=a\,\eta^{2}+2b\,\eta+c,$ (9.12)
with
$a=\frac{m^{2}+n^{2}}{m^{2}},\,\,b=\frac{n}{qm^{2}},\,\,c=\frac{1}{q^{2}m^{2}},\quad ac-b^{2}=\frac{1}{q^{2}m^{2}}$ (9.13)
Then, provided that $\alpha<\frac{1}{2}$, we obtain after the computation of an elementary integral
$\left|\overset{\circ}{I}_{n,m}\left(q^{2i\pi{\mathcal{E}}}h\right)\right|\leq\,C\,|m|^{-1}a^{-\alpha}(ac-b^{2})^{\alpha-\frac{1}{2}}=C\,(m^{2}+n^{2})^{-\alpha}q^{1-2\alpha}.$ (9.14)
We must improve the first factor to ensure the summability of the $(m,n)$-series, while improving slightly the second factor at the same time. Setting $k={\mathcal{F}}^{\mathrm{symp}}h$, one has with $k^{\prime}_{1}(x,\xi)=\frac{\partial k}{\partial x}$
$(2i\pi{\mathcal{E}}\,k)\left(\frac{n\eta}{m}+\frac{1}{qm},\,\eta\right)\\ =\left(1+\eta\frac{d}{d\eta}\right)\,\left[k\left(\frac{n\eta}{m}+\frac{1}{qm},\,\eta\right)\right]+\frac{1}{qm}\,k^{\prime}_{1}\left(\frac{n\eta}{m}+\frac{1}{qm},\,\eta\right)\\ =\frac{d}{d\eta}\,\left[\eta\,k\left(\frac{n\eta}{m}+\frac{1}{qm},\,\eta\right)\right]+\frac{1}{qm}\,k^{\prime}_{1}\left(\frac{n\eta}{m}+\frac{1}{qm},\,\eta\right).$ (9.15)
From (9.11), one obtains
$\overset{\circ}{I}_{n,m}\left((2i\pi{\mathcal{E}})\,q^{2i\pi{\mathcal{E}}}h\right)=\frac{1}{q\,|m|_{1}^{2}}\,\int_{-\infty}^{\infty}k^{\prime}_{1}\left(\frac{n\eta}{m}+\frac{1}{qm},\,\eta\right)\,d\eta.$ (9.16)
With respect to (9.11), we have gained a factor $\frac{1}{qm}$. Doing this integration by parts twice, we gain a factor $\frac{1}{q^{2}m^{2}}$. If substituting for the equation (9) the equation (9), we obtain in a similar way a gain by a factor $\frac{1}{q^{2}n^{2}}$. Hence, for $\alpha<\frac{1}{2}$,
$\left|\overset{\circ}{I}_{n,m}\left(q^{2i\pi{\mathcal{E}}}(2i\pi{\mathcal{E}})^{2}h\right)\right|\leq C\,(m^{2}+n^{2})^{-\alpha-1}q^{-1-2\alpha}.$ (9.17)
The $m,n$-series
$\big\langle\,{\mathfrak{B}}^{1},\,q^{2i\pi{\mathcal{E}}}\,h\big\rangle\,=\frac{1}{2}\sum_{m\equiv 0\,{\mathrm{mod}}\,N,\,n}I_{n,m}\left((\pi{\mathcal{E}})^{2}\,q^{2i\pi{\mathcal{E}}}h\right)$ (9.18)
is convergent if $\alpha>0$, and its sum is a ${\mathrm{O}}(q^{-1-2\alpha})$.
We have obtained that $q^{2i\pi{\mathcal{E}}}{\mathfrak{B}}^{j}$ remains in a bounded subset of ${\mathcal{S}}^{\prime}(\mathbb{R}^{2})$ for $q\geq 1$, As $q^{-2i\pi{\mathcal{E}}}{\mathfrak{B}}^{j}=q^{-2i\pi{\mathcal{E}}}{\mathcal{F}}^{\mathrm{symp}}\,{\mathfrak{B}}^{j}={\mathcal{F}}^{\mathrm{symp}}\left(q^{2i\pi{\mathcal{E}}}{\mathfrak{B}}^{j}\right)$, the proof of Proposition 9.1 is complete.
Theorem 9.2.
(Selberg’s conjecture) Given $M$, there is no cusp-form $\mathcal{N}$} for the group $\Gamma=\Gamma_{0}(M)$ such that $\Delta\,{\mathcal{N}}=\frac{1-\nu^{2}}{4}\,{\mathcal{N}}$ with ${\mathrm{Re\,}}(\nu^{2})>0$.
Proof.
The proof follows the same lines as that of our attempt of the Ramanujan case. But it is Proposition 9.1 that takes the place of Theorem 8.2, the family of powers of the Hecke operator $T_{p}^{\mathrm{dist}}$ being replaced by the family of operators $q^{2i\pi{\mathcal{E}}}$ with $q\in]0,\infty[$: this entails a considerable simplification.
In close analogy with what was done in Section 6, introduce the family of eigenvalues $\frac{1-\nu_{r}^{2}}{4},\,r=1,2,\dots$, corresponding to which there is at least one new Hecke eigenform ${\mathcal{N}}^{r}$: again, given $r$, we let $\left({\mathcal{N}}^{r,\iota}\right)_{\iota}$ be a complete orthogonal set of Hecke-normalized new eigenforms corresponding to this eigenvalue of $\Delta$ in $\)L^{2}(\Gamma\backslash\mathbb{H}$).
As a consequence of [3, p.226] and [3, p.244], one has
$\frac{(4\pi)^{j}}{\pi^{\frac{1}{2}}}\,\frac{\Gamma(\frac{1}{2}+j)}{\Gamma(\frac{1}{2})}\,U_{1}(\,\centerdot\,,\,j+\frac{1}{2})=\frac{1}{2}\sum_{r\geq 1,\iota}\frac{\Gamma\left(j-\frac{\nu_{r}}{2}\right)\Gamma\left(j+\frac{\nu_{r}}{2}\right)}{\|\,{\mathcal{N}}^{r,\iota}\,\|^{2}}\,{\mathcal{N}}^{r,\iota}+\dots,\\$ (9.19)
where the dots stand for a linear combination of Eisenstein series (relative to the various cups of the domain) and make up the continuous part of the spectral decomposition of the left-hand side.
The two Hecke distributions ${\mathfrak{N}}^{\pm r,\iota}$ being the lifts of ${\mathcal{N}}^{r,\iota}$ according to (9.2), define the distribution
$\left({\mathfrak{B}}^{j}\right)^{\mathrm{disc}}=\frac{1}{2}\sum_{r\neq 0,\,\iota}\frac{\Gamma\left(j-\frac{\nu_{r}}{2}\right)\Gamma\left(j+\frac{\nu_{r}}{2}\right)}{\|\,{\mathcal{N}}^{r,\iota}\,\|^{2}}\,{\mathfrak{N}}^{r,\iota}.$ (9.20)
It is invariant under ${\mathcal{F}}^{\mathrm{symp}}$ and its image under $\Theta$ coincides with the right-hand side of (9.19). One thus has
${\mathfrak{B}}^{j}=\left({\mathfrak{B}}^{j}\right)^{\mathrm{disc}}+\dots,$ (9.21)
where the dots stand for a linear combination of Eisenstein distributions ${\mathfrak{E}}_{i\lambda},\,\lambda\in\mathbb{R}$ or transforms thereof under the linear transformations of $\mathbb{R}^{2}$ associated to the cusps of the domain.
From Proposition 9.1 and the fact that $q^{2i\pi{\mathcal{E}}}{\mathfrak{E}}_{i\lambda}=q^{-i\lambda}{\mathfrak{E}}_{i\lambda}$, it follows, using also the analogue of Proposition 6.1, that the distributions $q^{2i\pi{\mathcal{E}}}\left({\mathfrak{B}}^{j}\right)^{\mathrm{disc}}$ make up a bounded set in ${\mathcal{S}}^{\prime}(\mathbb{R}^{2})$ for $q\in]0,\infty[$. The same is true if one replaces $\left({\mathfrak{B}}^{j}\right)^{\mathrm{disc}}$ by the part of (9.20) made from the lifts of all new Hecke eigenforms ${\mathcal{N}}^{r,\iota}$ corresponding to eigenvalues of $\Delta$ not less than $\frac{1}{4}$. We are left with a finite linear combination of terms $q^{2i\pi{\mathcal{E}}}{\mathfrak{N}}_{r,\iota}=q^{-\nu_{r}}{\mathfrak{N}}_{r,\iota}$, which remains in a bounded subset of ${\mathcal{S}}^{\prime}(\mathbb{R}^{2})$ for $q\in]0,\infty[$. Given $\nu_{1},\iota_{1}$, the result of testing the term $q^{2i\pi{\mathcal{E}}}{\mathfrak{N}}_{r_{1},\iota_{1}}$ on an appropriate function in ${\mathcal{S}}(\mathbb{R}^{2})$ is, just as well, bounded for $q\in]0,\infty[$ as a result of Lemma 8.1. Hence, $\nu_{1}$ is pure imaginary, and the eigenvalue $\frac{1-\nu_{1}^{2}}{4}$ is at least $\frac{1}{4}$. ∎